### v1

# Part I: Nonlinear Codes

Abstract
A de Bruijn arrays code is a set of $r\times s$ binary doubly-periodic arrays such that each binary $n\times m$ matrix is contained exactly once as a window in one of the arrays. Such a set of arrays can be viewed as a two-dimensional generalization of a perfect factor in the de Bruijn graph. Necessary conditions for the existence of such arrays are given. Several direct constructions and recursive constructions for such arrays are given. A framework for a theory of two-dimensional feedback shift register which is akin to (one-dimensional) feedback shift registers is suggested.
I Introduction
Generalizations of one-dimensional sequences and codes to higher dimensions is quite natural from both theoretical and practical point of view. Such generalizations were considered for various structures such as error-correcting codes [1, 47], burst-correcting codes [2, 3, 15, 16], constrained codes [50, 54], and de Bruijn sequences [8, 34, 42]. This paper considers a generalization of one-dimensional sequences with a window property to two-dimensional arrays with a window property. For simplicity in this paper only binary arrays and usually binary sequences are considered, although some of the results can be generalized to any alphabet size and some sequences over larger alphabets can be used to construct binary arrays.
The de Bruijn graph $G_{\sigma,n}=(V_{n},E_{n})$ ($G_{n}$ when $\sigma=2$) is a directed graph, where the set of vertices $V_{n}$ has $\sigma^{n}$ vertices represented by the set of $\sigma^{n}$ words of length $n$ over an alphabet of size $\sigma$. The set of edges $E_{n}$ consists of $\sigma^{n+1}$ directed edges represented by the $\sigma^{n+1}$ words of length $n+1$. The edge $(x_{0},x_{1},\ldots,x_{n-1},x_{n})$ is directed from the vertex $(x_{0},x_{1},\ldots,x_{n-1})$ to the vertex $(x_{1},\ldots,x_{n-1},x_{n})$. A span $n$ de Bruijn sequence is a cyclic sequence of length $\sigma^{n}$ in which each $n$-tuple is contained in exactly one window of $n$ consecutive digits. This cyclic sequence is equivalent to an Eulerian cycle in $G_{\sigma,n-1}$ and a Hamiltonian cycle in $G_{\sigma,n}$. A span $n$ shortened de Bruijn sequence is a sequence of length $\sigma^{n}-1$ in which each nonzero $n$-tuple is contained in exactly one window of $n$ consecutive digits. The following definition generalizes the definition of a de Bruijn sequence to a two-dimensional array.
Definition 1
A de Bruijn array (or a perfect map) is an $r\times s$ doubly-periodic array, such that each $n\times m$ binary matrix appears exactly once as a window in the array. Such an array will be referred to as $(r,s;n,m)$-DBA.
Perfect maps were first presented by Reed and Stewart [46] and later in Gordon [21] and Clapham [6]. A very simple construction based on a span $n$ de Bruijn sequence and a de Bruijn sequence over an alphabet of size $2^{n}$ was presented by Ma [31]. This construction will be modified later in one of our constructions. Fan, Fan, Ma, and Siu [17] gave a two-dimensional graphical representation for perfect maps. They also presented two recursive constructions for such de Bruijn arrays. An important construction for perfect maps and another graphical representation was given in [8]. Finally, Paterson [42] proved that the necessary conditions for the existence of de Bruijn arrays are also sufficient. These arrays have found a variety of applications. They were used in pattern recognition for structured light systems as described in Geng [19], Lin, Nie, and Song [29], Morano, Ozturk, Conn, Dubin, Zietz, and Nissanov [40], Salvi, Fernandez, Pribanic, and Llado [48], and Salvi, Pagès, and Batlle [49]. They are also used in transferring planar surface into a sensitive touch screen display, see Dai and Chung [7], in camera localization as described by Szentandrasi, Zachariás̆, Havel, Herout, Dubska, and Kajan [53], and in one-shot shape acquisition, e.g., Pagès, Salvi, Collewet, and Forest [41]. Finally, they can be applied in surface measurements as described in Kiyasu, Hoshino, Yano, and Fujimura [26] and in Spoelder, Vos, Petriu, and Groen [52], and also in coded aperture imaging as was presented for example in Gottesman and Fenimore [22].
The are of de Bruijn arrays (which were generally called perfect maps) were extensively covered in the literature. Some more work on this topic and related structures was done by [23, 24, 25, 30, 32, 33, 34, 35, 36, 37]
The goal of the current paper is to extend the results on de Bruijn arrays into de Bruijn arrays codes defined as follows.
Definition 2
A de Bruijn arrays code is a set of $r\times s$ doubly-periodic arrays, such that each nonzero $n\times m$ matrix appears exactly once as a window in one of the arrays. Such a set of arrays will be referred to as $(r,s;n,m)$-DBAC.
It is straightforward to verify the necessary conditions for the existence of an $(r,s;n,m)$-DBAC.
Lemma 1
If $\mathbb{C}$ is an $(r,s;n,m)$-DBAC of size $\Delta$, then
  1. 1.
    $r>n$ or $r=n=1$,
  2. 2.
    $s>m$ or $s=m=1$.
  3. 3.
    $\Delta rs=2^{nm}$.
Lemma 1 implies that $r$, $s$, and $\Delta$, are powers of two. Henceforth, we use $2^{k}$ instead of $r$ and $2^{t}$ instead of $s$. One of the goals in the research on DBACs is to prove that the necessary conditions of Lemma 1 are sufficient. In the discussion which is given in this paper it will be also shown how the definition of de Bruijn arrays codes generalizes some one-dimensional definitions.
The rest of this paper is organized as follows. In Section II we present the basic definitions and theory for the arrays which will be considered in this work. The section consists of five subsections. To understand the theory we have to understand some basic concepts associated with shift registers and their sequences. These will be described in Section II-A. Some state diagrams of these registers are the one-dimensional analogs of the de Bruijn arrays codes. Two well-known operators defined for the one-dimensional theory of shift registers will play an important role in our exposition, the shift operator ${\bf E}$, and the ${\bf D}$-morphism. The shift operator ${\bf E}$ will be discussed in Section II-B and the emphasis will be on its role in defining sequences which are generated by linear feedback shift registers. It will be also important in the definition of linear complexity of sequences which will be used for some of the constructions of the two-dimensional arrays. The operator ${\bf D}$ will be discussed in Section II-C. This operator defines a homomorphism between de Bruijn graphs of different orders. It will be used in the recursive constructions of de Bruijn arrays and it is strongly connected to the shift operator ${\bf E}$. The structures that will be heavily used in both the direct constructions and the recursive constructions are the perfect factors which will be discussed in Section II-D. Perfect factors were used in the constructions of de Bruijn arrays and will be used in the constructions of de Bruijn arrays codes. They will be used in the direct constructions and in the recursive constructions. In the recursive constructions they will be used as columns of the arrays and also as sequences which define how the constructions are applied with the inverse of the operator ${\bf D}$. The one-dimensional sequences are associated with the de Bruijn graph and for the two-dimensional arrays two different graph representations will be presented in Section II-E. In particular we will examine how one-dimensional nonsingular feedback shift registers are generalized into two-dimensions. Many constructions will be presented in the paper, some of these constructions are direct constructions and some of them are recursive constructions. Section III is devoted to five direct constructions. One construction is obtained by ordering the sequences of a perfect factor. In a second construction the sequences of a perfect factor are ordered in a way that all the possible shifts between two sequences are obtained. A third construction generalizes the construction of Ma [31] using a perfect factor over a large alphabet and a binary de Bruijn sequence. The last two direct constructions are two-dimensional generalizations of two of the most celebrated nonsingular shift registers. These generalizations should be done carefully to obtain a state diagram with cycles of the same length. Five recursive constructions will be presented in Section IV. All these constructions are generalizations of constructions of de Bruijn arrays presented by Fan, Fan, Ma, and Siu [17]. Perfect factors have to be used carefully for these constructions and the chosen arrays on which the constructions are applied should be also chosen in an appropriate way. In Section V another technique to generate $(2^{k},2^{t};n,m)$-DBACs by joining codewords (cycles) of an $(2^{k},2^{t^{\prime}};n,m)$-DBAC with shorter codewords, i.e., $t^{\prime}<t$, is presented. This technique can be applied if all the codewords are cycles of the same length in the same state diagram of the two-dimensional generalization de Bruijn graph. Section VI includes conclusion and future work.
MacWilliams and Sloane [32] gave a construction for another type of arrays defined as follows.
Definition 3
A pseudo-random array ${\cal A}$ is an $r\times s$ doubly-periodic array such that each $n\times m$ nonzero matrix is contained exactly once as a window in the array. Moreover, if ${\cal A}$ is such an array, then ${\cal A}+{\cal A}^{\prime}$, where ${\cal A}^{\prime}$ is a nontrivial shift of ${\cal A}$ and the addition is performed bit-by-bit, is another nontrivial shift of ${\cal A}$. This is a shift-and-add-property. Such an array will be called an $(r,s;n,m)$-PRA.
Definition 4
A shortened perfect map (or a shortened de Bruijn array) is an $r\times s$ binary doubly-periodic array, where ${rs=2^{nm}-1}$, such that each nonzero $n\times m$ matrix appears exactly once as a window in the array. Such an array will be called an $(r,s;n,m)$-SDBA.
The definition of pseudo-random arrays is generalized to pseudo-random arrays codes and it will be considered in the follow-up paper [5]. They were also considered in a preliminary version that discussed these array codes [11]. In the current paper the constructions of DBACs will be based on combinatorial techniques and the constructed codes will be nonlinear. In the follow-up paper [5] the constructions of pseudo-random arrays codes will be based on algebraic techniques and the constructed codes will be linear.
II Preliminaries
This section is devoted to the basic concepts which will be used in our constructions of DBACs and also in explaining how the one-dimensional setup is generalized to the two-dimenstional setup.
II-A Shift-register sequences
A feedback shift register of order $n$ (an FSRn in short) has $2^{n}$ states, represented by the set of the $2^{n}$ binary words of length $n$. The register has $n$ cells, which are binary storage elements, where each cell stores at each stage one of the bits of the current state $x=(x_{1},x_{2},\ldots,x_{n})$. An FSRn is depicted in Fig. 1.
Figure 1: Feedback shift register of order $n$.
Shift registers and their sequences were extensively studied and two comprehensive books on them are [10, 20]. The known material in this section is presented in these books. Given the word $(x_{1},x_{2},\ldots,x_{n})$ which is a state in the FSRn, $x_{i}$ is stored in the $i$-th cell of the FSRn. The $n$ cells are connected to another logic element which computes a Boolean feedback function $f(x_{1},x_{2},\ldots,x_{n})$. At periodic intervals which are controlled by a master clock, $x_{2}$ is transferred to $x_{1}$, $x_{3}$ to $x_{2}$, and so on until $x_{n}$ is transferred to $x_{n-1}$. The value of the feedback function is transferred to $x_{n}$ and hence it is common to denote $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. The register starts to work with an initial state $(a_{1},a_{2},\ldots,a_{n})$, where $a_{i}$, $1\leq i\leq n$, is the initial value stored in the $i$-th cell. The feedback function $f$ is a Boolean function and hence it can be also represented by a truth table. There are $2^{n}$ possible distinct values for $x_{1},x_{2},\ldots,x_{n}$, i.e., there are $2^{n}$ distinct states and each one can have a value of either 0 or 1. Hence, there are $2^{2^{n}}$ different FSRn, but not all of these functions are of interest.
Each FSRn has a state diagram which is a graph with $2^{n}$ vertices (the states of the FSRn). Given an FSRn with a feedback function $f(x_{1},x_{2},\ldots,x_{n})$, the vertex $(x_{1},x_{2},\ldots,x_{n})$ in the associated state diagram has an edge to the vertex $(x_{2},\ldots,x_{n},x_{n+1})$ if $x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})$. This implies that the state diagram with $2^{n}$ states (vertices) has exactly $2^{n}$ edges. A nonsingular feedback shift register has a feedback function whose state diagram contains only cycles. Similarly, a factor in a graph is a set of vertex-disjoint cycles that contain all the vertices in the graph. A factor in $G_{n}$ is associated with a state diagram of a nonsingular feedback shift register. These are the only FSRns which are of interest and hence in the rest of the paper all FSRns will always be nonsingular. We can easily verify if a feedback shift register is nonsingular by the following lemma.
Lemma 2
An $\textup{FSR}_{n}$ is nonsingular if and only if for each $x_{i}\in\{0,1\}$, $2\leq i\leq n$, we have
$f(0,x_{2},\ldots,x_{n})\neq f(1,x_{2},\ldots,x_{n})$
Corollary 1
An $\textup{FSR}_{n}$ is nonsingular if and only if for each $x_{i}\in\{0,1\}$, $2\leq i\leq n$, we have
$f(x_{1},x_{2},\ldots,x_{n})=x_{1}+g(x_{2},\ldots,x_{n})~{},$
where $g(x_{2},\ldots,x_{n})$ is any Boolean function with $n-1$ variables.
A nonsingular linear feedback shift register of order $n$ (an LFSRn in short) over $\mathbb{F}_{q}$ is an FSRn whose feedback function $f$ is linear, i.e.,
$x_{n+1}=f(x_{1},x_{2},\ldots,x_{n})=x_{1}+\sum_{i=2}^{n}c_{i}x_{i},~{}~{}c_{i}\in\mathbb{F}_{2},~{}~{}1\leq i\leq n~{}.$
The class of linear shift registers is the most important one, but most of the shift registers are nonlinear. While, by Corollary 1 there are $2^{2^{n-1}}$ nonsingular feedback shift registers of order $n$, the definition of LFSRn implies that there are only $2^{n-1}$ nonsingular LFSRns. Some functions of nonlinear shift register are very similar to the functions of the linear ones and these will be of special interest. Four feedback functions will be defined now.
The first one is called the PCRn for the pure cycling register of order $n$, and its linear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=x_{1}~{}.$
The second FSRn is called the CCRn for the complemented cycling register of order $n$, and its nonlinear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=x_{1}+1~{}.$
The third FSRn, which will be considered, is called the PSRn for the pure summing register of order $n$. Its linear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{i=1}^{n}x_{i}~{}.$
The fourth FSRn is called the CSRn for the complemented summing register of order $n$, and its nonlinear feedback function is
$f(x_{1},x_{2},\ldots,x_{n})=\sum_{i=1}^{n}x_{i}+1~{}.$
The length of the cycles in the state diagrams of these shift-registers are of special interest. The length of the cycle in the state diagram is the period of the associated sequence. The length of the sequence is the number of symbols written in the sequence. Hence, the length of the sequence $S=[s_{0}s_{1}\cdots s_{n-1}]$ is $n$. The period $\pi(S)$ of the cyclic sequence ${\cal S}=[s_{0}s_{1}\cdots s_{n-1}]$, is the least positive integer $\pi$ such that $s_{i}=s_{\pi+i}$, for each $10\leq i\leq n-$ 1, where indices are taken modulo $n$, when the sequence is cyclic. For a cyclic sequence ${\cal S}=[s_{0}s_{1}\cdots s_{n-1}]$ any integer $\pi$, $1\leq\pi\leq n$, such that $\pi$ divides $n$ and $s_{i}=s_{\pi+i}$ for $10\leq i\leq n-$ 1 is a period of $S$, but the period is the smallest such integer. Two cyclic sequences are said to be equivalent if one is a cyclic shift of the other. Such sequences are usually considered to be the same unless the starting point of the cyclic sequence is determined. Similar definitions for the length and the period are applied on arrays, The length and the period of sequences generated from shift registers were considered in [10, 20]. These comprehensive books, on shift-register sequences, the de Bruijn graph and its sequences, also consider the length of cycles in any state diagram of linear shift register. A linear shift register can be represented and analyzed by polynomials and it will be discussed also in Section II-B. A cycle of the state diagram can be represented by a sequence whose length equal to the length of the cycle. The sequence is generated from the first bits of the vertices in the cycle keeping their order as the order of the vertices in the cycle.
Example 1
The cyclic sequence
$[010010010]$
is of length 9, but its period is 3. It also has periods 6 and 9.
The cyclic sequence
$[00101]$
is of length 5 and its period is 5.
The two sequences
$[0011101111001]~{}~{}~{}\text{and}~{}~{}~{}[1011110010011]$
are equivalent sequences as one is a cyclic shift of the other, i.e.,
$[0011101111001]\simeq[1011110010011].$
$\blacksquare$
The companion of a vertex (state) $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the vertex $x^{\prime}=(x_{1},x_{2},\ldots,x_{n-1},\bar{x}_{n})$, where $\bar{b}$ is the binary complement of $x$. The complement of a state $x=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ is the state $\bar{x}=(\bar{x}_{1},\bar{x}_{2},\ldots,\bar{x}_{n-1},\bar{x}_{n})$. A cycle (sequence) ${\cal S}=[s_{1},s_{2},\ldots,s_{k}]$ is called self-dual if the complement of ${\cal S}$, ${\bar{{\cal S}}=[\bar{s}_{1},\bar{s}_{2},\ldots,\bar{s}_{k}]}$ is equivalent to ${\cal S}$. A sequence ${\cal S}$ is self-dual if and only if it can has the form ${\cal S}=[X~{}\bar{X}]$.
Two cycles ${\cal C}_{1}$ and ${\cal C}_{2}$ with a state $x$ on ${\cal C}_{1}$ and its companion $x^{\prime}$ on ${\cal C}_{2}$ are merged into one cycle ${\cal C}$ when the predecessors of $x$ and $x^{\prime}$, respectively, $z$ on ${\cal C}_{1}$ and $y$ on ${\cal C}_{2}$, respectively, are interchanged. This operation is called the merge-or-split method since it either merges two cycles into one cycle or it splits one cycle into two cycles. The two possible scenarios of the merge-or-split method are depicted in Fig. 2.
Figure 2: The merge-or-split method.
II-B The shift operator ${\bf E}$
A very important operator with whom many of the properties of feedback shift registers can be represented is the shift operator ${\bf E}$. This operator is shifting the whole sequence to the left. For a sequence ${\cal S}=s_{0}s_{1}~{}\cdots~{}s_{k-1}$ of length $k$, the shift operator ${\bf E}s_{i}=s_{i+1}$, where $10\leq i<k-$ 1 and if the sequence is cyclic then also ${\bf E}s_{k-1}=s_{0}$. When the operator is applied on the whole cyclic sequence ${\cal S}$ we have that
${\bf E}[s_{0}s_{1}~{}\cdots~{}s_{k-1}]=[s_{1}~{}\cdots~{}s_{k-1}s_{0}]~{}.$
When the operator is applied on an acyclic sequence ${\cal S}$ we have that
${\bf E}(s_{0}s_{1}~{}\cdots~{}s_{k-1})=(s_{1}~{}\cdots~{}s_{k-1})~{}.$
Any periodic sequence satisfies a linear recursion
$s_{i+m}=a_{1}s_{i+m-1}+\cdots+a_{m-1}s_{i+1}+a_{m}s_{i}~{},~{}i\geq 0,$ (1)
of order $m\leq k$, where $a_{j}\in\{0,1\}$ and all computations are done modulo 2. In terms of the shift operator ${\bf E}$ the linear recursion (1) takes the form
${\bf E}^{m}s_{i}=\sum_{j=1}^{m}a_{j}{\bf E}^{m-j}s_{i}~{}~{}\text{or}~{}~{}\left({\bf E}^{m}+\sum_{j=1}^{m}a_{j}{\bf E}^{m-j}\right)s_{i}=0~{},~{}~{}~{}~{}i\geq 0.$ (2)
If $m$ is the smallest such integer, it implies that ${a_{m}\neq 0}$. If we define
$f({\bf E})\triangleq{\bf E}^{m}+\sum_{j=1}^{m}a_{j}{\bf E}^{m-j}~{},$
then we have $f({\bf E})s_{i}=0$ for each $i\geq 0$. This implies that (2) takes the form
$f({\bf E})s_{i}=\left({\bf E}^{m}+\sum_{j=0}^{m-1}c_{j}{\bf E}^{j}\right)s_{i}=0~{},$
for each $i\geq 0$. In this case it is said that the polynomial $f({\bf E})$ generates the sequence ${\cal S}$ (or ${\cal S}$ is generated by $f({\bf E})$).
Henceforth, we denote by ${\bf 0}$ any sequence (or a matrix) of any size which contains only zeros. The linear complexity $c({\cal S})$ of ${\cal S}$ is defined as the least $m$ for which (1) holds. Clearly ${c({\cal S})\leq k}$, for any sequence with period $k$, since $s_{i+k}=s_{i}$ by definition.
The sequences that are of interest in our exposition have length which is a power of 2. These sequences are generated by the polynomial $({\bf E}+1)^{k}$ and for each sequence ${\cal S}$ generated by this polynomial we have that $({\bf E}+{\bf 1})^{k}{\cal S}={\bf 0}$. The following properties characterize these sequences [10].
Lemma 3
A sequence ${\cal S}$ of length $2^{n}$ has linear complexity $c$ if and only if $({\bf E}+{\bf 1})^{c-1}{\cal S}=1^{2^{n}}$.
Lemma 4
There are $2^{n-\lfloor\log n\rfloor-1}$ sequences which are generated by the polynomial $({\bf E}+{\bf 1})^{n+1}$ and are not generated by the polynomial $({\bf E}+{\bf 1})^{n}$. All these sequences have period $2^{\lfloor\log n\rfloor+1}$ and linear complexity ${n+1}$. Each $n$-tuple is contained exactly once in one of these sequences.
We define the weight of a word $(x_{1},x_{2},\ldots,x_{n})$ as the number of nonzero entries in the word. Similarly, we define the weight of a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ as the number of nonzero entries in the sequence. Note, that the weight of the word $(011)$ is two, of the sequence $[011]$ is also two, but of the sequence $[011011]$ is four, although its period is three and the weight of one period is two. Some other properties of sequences can be characterized as follows.
Lemma 5
For $n\geq 1$, a sequence ${\cal S}$ of length $2^{n}$ has linear complexity $c>0$ if and only if the linear complexity of the sequence $({\bf E}+{\bf 1}){\cal S}$ is $c-1$.
Lemma 6
For $n\geq 1$, a sequence ${\cal S}$ of length $2^{n}$ has linear complexity $2^{n}$ if and only if its weight is odd.
Lemma 7
For $n\geq 1$, a sequence ${\cal S}$ of length $2^{n+1}$ has linear complexity $12^{n}+$ 1 if and only if ${\cal S}$ is a self-dual sequence.
Finally, we consider the linear complexity of de Bruijn sequences [4, 9, 12, 13, 18].
Theorem 1
For $n\geq 3$, if ${\cal S}$ is a span $n$ de Bruijn sequence, then $12^{n-1}+n\leq c({\cal S})\leq 2^{n}-$ 1 and $c({\cal S})\neq 2^{n-1}+n+1$. There exists a span $n$ de Bruijn sequence ${\cal S}$ for which $c({\cal S})=2^{n-1}+n$ and a span $n$ de Bruijn sequence ${\cal S}$ for which $c({\cal S})=2^{n}-1$.
Conjecture 1
If $n\geq 3$ and $c$ is an integer such that $12^{n-1}+n\leq c\leq 2^{n}-$ 1 and $c\neq 2^{n-1}+n+1$, then there exists a span $n$ de Bruijn sequence such that $c=c({\cal S})$.
II-C The ${\bf D}$-morphism
A key operator in the recursive constructions which follow is the ${\bf D}$-morphism which was defined by Lempel [27]. The operator ${\bf D}$ is a 2-to-1 mapping from $\mathbb{F}_{2}^{n}$ on to $\mathbb{F}_{2}^{n-1}$, where $n\geq 2$. For a binary word $(x_{1},x_{2},\ldots,x_{n})$ we have
${\bf D}(x_{1},x_{2},\ldots,x_{n})=(x_{1}+x_{2},x_{2}+x_{3},\ldots,x_{n-1}+x_{n})=({\bf E}+{\bf 1})(x_{1},x_{2},\ldots,x_{n}).$
By this definition we have that ${\bf D}x={\bf D}y$ if and only if $y=x$ or $y=\bar{x}$. The operator ${\bf D}$ can be apply on cyclic sequences (cycles). For a cyclic sequence $[x_{1},x_{2},\ldots,x_{n}]$ we have
${\bf D}[x_{1},x_{2},\ldots,x_{n}]=[x_{1}+x_{2},x_{2}+x_{3},\ldots,x_{n-1}+x_{n},x_{n}+x_{1}]=({\bf E}+{\bf 1})[x_{1},x_{2},\ldots,x_{n}].$
The operator ${\bf D}$ has an inverse ${\bf D}^{-1}$. For a sequence ${\cal S}=[x_{1},x_{2},\ldots,x_{n}]$ of even weight, ${\bf D}^{-1}{\cal S}$ is a set that contains two complementary sequences of length $n$,
${\bf D}_{0}^{-1}{\cal S}=[0,x_{1},x_{1}+x_{2}+x_{3},\cdots,\sum_{i=1}^{n-1}x_{i}]$
and
${\bf D}_{1}^{-1}{\cal S}=[1,1+x_{1},1+x_{1}+x_{2}+x_{3},\cdots,1+\sum_{i=1}^{n-1}x_{i}].$
For a sequence ${\cal S}=[x_{1},x_{2},\ldots,x_{n}]$ of odd weight, ${\bf D}^{-1}{\cal S}$ contains one self-dual sequence that can be written as
${\bf D}_{0}^{-1}{\cal S}=[0,x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{n-1}x_{i},1,1+x_{1},1+x_{1}+x_{2},\ldots,1+\sum_{i=1}^{n-1}x_{i}]$
or as
${\bf D}_{1}^{-1}{\cal S}=[1,1+x_{1},1+x_{1}+x_{2},\ldots,1+\sum_{i=1}^{n-1}x_{i},0,x_{1},x_{1}+x_{2},\ldots,\sum_{i=1}^{n-1}x_{i}].$
The definition of ${\bf D}$ implies that ${\bf D}={\bf E}+{\bf 1}$ and hence some of the results concerning the shift operator ${\bf E}$ associated with sequences of length $2^{n}$ can be given in terms of the operator ${\bf D}$, e.g.,
Lemma 8
For a nonzero sequence ${\cal S}$ of length $2^{n}$ we have $c({\cal S})=c({\bf D}{\cal S})+1$.
Finally, we can apply the inverse operator ${\bf D}^{-1}$ on a set of sequences ${\cal T}$ as follows:
${\bf D}^{-1}{\cal T}\triangleq\{{\bf D}^{-1}{\cal S}~{}:~{}{\cal S}\in{\cal T}\}$
Lemma 9
If ${\cal F}$ is a factor in $G_{n}$, then ${\bf D}^{-1}{\cal F}$ is a factor in $G_{n+1}$.
The operator ${\bf D}$ and its inverse ${\bf D}^{-1}$ can be applied several consecutive times on a sequence ${\cal S}$ and the derived sequences. It is defined for $k\geq 2$ as follows.
${\bf D}^{k}{\cal S}=({\bf E}+{\bf 1})^{k}{\cal S}=({\bf E}+{\bf 1})({\bf E}+{\bf 1})^{k-1}{\cal S}$
and
${\bf D}^{-k}{\cal S}={\bf D}^{-1}({\bf D}^{-k+1}){\cal S}$
The equality ${\bf D}={\bf E}+{\bf 1}$ and the properties the operator ${\bf D}$, its inverse ${\bf D}^{-1}$, and the linear complexity imply the following consequences.
Lemma 10
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length $2^{k}$, , with linear complexity smaller than $2^{k}$, in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $2\Delta$ vertex-disjoint cycles of length $2^{k}$ in $G_{n+1}$.
Lemma 11
If ${\cal F}$ is a set of $\Delta$ vertex-disjoint cycles of length $2^{k}$, with linear complexity $2^{k}$, in $G_{n}$, then ${\cal D}^{-1}{\cal F}$ contains $\Delta$ vertex-disjoint self-dual cycles of length $2^{k+1}$ in $G_{n+1}$.
II-D Perfect factors
The most important factors in $G_{n}$ which will play a pivotal role in our exposition are the perfect factors.
Definition 5
[…]
The following theorem on the existence of perfect factors was proved in [8].
Theorem 2
A perfect factor PF$(n,k)$ exists if and only if $k\leq n<2^{k}$, i.e., $n<2^{k}\leq 2^{n}$.
The theorem was extended for larger alphabet by Paterson [43]. The following case of his result is required for Theorem 7.
Theorem 3
[…]
For a given set of parameters there could be different perfect factors. We will be interested in perfect factors in which the weights of all the sequences are either even or odd. Moreover, in these perfect factors all the sequences have the same linear complexity. Both types of factors will be important in our exposition. For sequences of even weight we can characterize all the parameters of perfect factors.
Theorem 4
  1. 1.
    If $k\leq n<2^{k}-1$, then there exists a PF$(n,k)$ for which all the sequences are of even weight.
  2. 2.
    If $n=2^{k}-1$ then there exists a unique PF$(n,k)$ and all the sequences in this factor are of odd weight.
Proof:
  1. 1.
    If $12^{k-1}\leq n<2^{k}-$ 1, then by Lemma 4 all the sequences of length $2^{k}$ which are generated by the polynomial $({\bf E}+{\bf 1})^{n+1}$ and are not generated by the polynomial $({\bf E}+{\bf 1})^{n}$ form a PF$(n,k)$. By Lemma 4 all these sequences have linear complexity $n+1$, where $2^{k-1}<n+1<2^{k}$ and hence by Lemma 6 all these sequences have even weight.
    If $k\leq n<2^{k-1}$, then consider a span $k$ de Bruijn sequence ${\cal S}$ with linear complexity $2^{k-1}+k$ whose existence is implied by Theorem 1. By Lemmas 810, and 11 we have that ${\bf D}^{-r}{\cal S}$ for any $r$ such that $0\leq r<2^{k-1}-k$ contains $2^{r}$ sequences of length $2^{k}$ in $G_{k+r}$, i.e., $2^{n-k}$ sequences of length $2^{k}$ in $G_{n}$ which is a PF$(n,k)$ for any given $n$ such that $k\leq n<2^{k-1}$. By Lemma 8 all the generated sequences have linear complexity $2^{k-1}+k+r<2^{k}$ and hence by Lemma 6 all these sequences have even weight.
  2. 2.
    For $n=2^{k}-1$ let ${\cal F}$ be perfect factor PF$(n,k)$ with sequences of length $n+1=2^{k}$. One of these sequences contains the all-zero word of length $n$. Clearly, the sequence which contains the all-zero word of length $n+1$ is $[0^{n}1]$. This sequence contains all words of length $n$ and weight one. Since each cyclic sequence of length $n+1$ and weight two contains exactly two words of weight one (by omitting a one from the sequence), it follows that ${\cal F}$ does not contain sequences of weight two and hence all words of weight two are contained in sequences with weight three. This implies that ${\cal F}$ does not contain sequences of weight four and so on by induction. Thus, there is a unique factor whose sequences have odd weight. This factor is the state diagram of the CSRn.
Theorem 4 implies that whenever a perfect factor PF$(n,k)$ exists, except for $n=2^{k}-1$, i.e., when $k\leq n<2^{k}-1$ by Theorem 2, there exists such a perfect factor for which all its sequences have even weight. The situation is more complicated for odd weight. PF$(n,n)$ contains exactly one sequence of length $2^{n}$ in $G_{n}$, i.e., a span $n$ de Bruijn sequence. Hence, its weight is always even. For all the other parameters a perfect factor in which all the sequences have odd weight can exist. Unfortunately, whether such perfect factors exist is an open problem. By Theorem 4, if $n=2^{k}-1$, then all the sequences in a PF$(n,k)$ have odd weight. For the other parameters we can use the same technique used in the proof of Theorem 4 and this is left for the reader. In other words we can start with a span $k$ de Bruijn sequence whose linear complexity is $2^{k-1}+\delta$, where $k\leq\delta\leq 2^{k-1}-k-1$ and $\delta\neq k+1$, and apply the inverse operator ${\bf D}^{-r}$, $r>0$, until we reach sequences with linear complexity $2^{n}$, where $n\geq k$. The obtained sequences will be of odd weight by Lemma 6. Other than these two techniques we do not know how to construct perfect factors whose sequences have only odd weights.
It is important to note that in all the constructions that were mentioned, if a sequence ${\cal S}$ is a sequence in the perfect factor ${\cal F}$, then $\bar{{\cal S}}$ is also a sequence in ${\cal F}$, unless ${\cal F}$ contains exactly one sequence. By Lemma 7 all these sequences of length $2^{k}$ are self-dual if and only if they are of linear complexity $12^{k-1}+$ 1. But, it should be noted that other perfect factors without these properties can be also constructed.
Example 2
When $k=4$ the length of the sequences is $162^{4}=$ 16 and there exists a PF$(n,4)$ with sequences of even weight for each $n$, $4\leq n\leq 14$. There exists a PF$(n,4)$ with sequences of odd weight for ${n\in\{5,6,8,12,15\}}$. For $n\in\{7,9,10,11,13,14\}$ we do not know whether such a perfect factor exists.
When $k=5$ the length of the sequences is $322^{5}=$ 32 and there exists a PF$(n,5)$ with sequences of even weight for each $n$, $5\leq n\leq 30$. There exists a PF$(n,5)$ with sequences of odd weight for $6\leq n\leq 14$ and $n\in\{16,21,22,24,28,31\}$. For $n\in\{15,17,18,19,20,23,25,26,27,29,30\}$ we do not know whether such a perfect factor exists.
Finally, one type of PF$(n,n-1)$ is of special interest. In this type of a perfect factor there are two sequences ${\cal S}$ and $\bar{{\cal S}}$. The sequence which contains the all-zero $n$-tuple in such a factor will be called a span $n$ half de Bruijn sequence. In such a sequence either $x$ or $\bar{x}$ is contained in ${\cal S}$, for any given $n$-tuple $x$.
II-E Graphical presentation of two-dimensional arrays
One-dimensional sequences are viewed as cycles in the de Bruijn graph. We aim to view the two-dimensional arrays that will be constructed in our work also as cycles in graphs which will be defined now.
A first interesting and important graph was presented by Fan, Fan, Ma, and Siu [17]. The graph $G_{n,m}(V,E_{1},E_{2})$ has a set of vertices $V$ which consists of all the $2^{nm}$ binary $n\times m$ matrices. There are two sets of edges in the graph, $E_{1}$ and $E_{2}$. The edges in $E_{1}$ are vertical edges and the edges in $E_{2}$ are horizontal edges. Two vertices ($n\times m$ matrices) $v$ and $u$ of $V$ are connected by an edge of $E_{1}$ from $u$ to $v$ if the $(n-1)\times m$ sub-matrix induced by the last $n-1$ rows of $u$ equals to the $(n-1)\times m$ sub-matrix induced by the first $n-1$ rows of $v$. Two vertices ($n\times m$ matrices) $v$ and $u$ of $V$ are connected by an edge of $E_{2}$ from $u$ to $v$ if the $n\times(m-1)$ sub-matrix induced by the last $m-1$ columns of $u$ equals to the $n\times(m-1)$ sub-matrix induced by the first $m-1$ columns of $v$.
A set of $2^{nm-k-t}$ doubly-periodic matrices of size $2^{k}\times 2^{t}$ which is an $(2^{k},2^{t};n,m)$-DBAC $\mathbb{C}$ is equivalent to a set of $2^{nm-k-t}$ disjoint components in the graph $G_{n,m}(V,E_{1},E_{2})$, where each component, associated with a $2^{k}\times 2^{t}$ matrix ${\cal C}\in\mathbb{C}$, contains $2^{k+t}$ vertices which are the $n\times m$ matrices induced by the $2^{k+t}$ windows of size $n\times m$ in ${\cal C}$. These vertices and their connecting edges in one component contain $2^{t}$ vertex-disjoint cycles of length $2^{k}$ with edges from $E_{1}$ and $2^{k}$ vertex disjoint cycles of length $2^{t}$ with edges from $E_{2}$. In other words, they form a factor with edges from $E_{1}$ and a factor with edges from $E_{2}$ in the subgraph induced by the vertices of the component induced by ${\cal C}$. Moreover, it is required that a cycle of length $2^{k}$ with edges from $E_{1}$ is associated with the $2^{k}$ $n\times m$ matrices obtained from a $2^{k}\times m$ sub-matrix of ${\cal C}$. A cycle of length $2^{t}$ with edges from $E_{2}$ is associated with the $2^{t}$ $n\times m$ matrices obtained from an $n\times 2^{t}$ sub-matrix of ${\cal C}$.
The second graph for the representation of the two-dimensional codes was defined before in [8]. This graph is appropriate to codes in which all the columns of the matrices that form the codewords are sequences of some given perfect factor. For this purpose, let ${\cal F}$ be a PF$(n,k)$ and define a graph $G_{{\cal F},m}$ whose vertices are $2^{k}\times m$ matrices. These matrices are periodic only vertically. Each such a matrix $X$ can be represented as
$X=(X_{i_{0}},{\bf E}^{j_{1}}X_{i_{1}},\ldots,{\bf E}^{j_{m-1}}X_{i_{m-1}}),$
where $X_{i_{\ell}}$, $10\leq\ell\leq m-$ 1, is a cycle in ${\cal F}$ represented as a column vector and ${\bf E}^{j_{\ell}}X_{i_{\ell}}$ is a cyclic shift up of $X_{i_{\ell}}$ by $j_{\ell}$ positions compared to its zero state, i.e., the column vector of ${\bf E}^{j_{\ell}}X_{i_{\ell}}$. The first cycle $X_{i_{0}}$ is taken in its zero shift. An edge in the graph is a $2^{k}\times(m+1)$ matrix which is periodic only vertically. Such a matrix
$Y=(Y_{i_{0}},{\bf E}^{j_{1}}Y_{i_{1}},\ldots,{\bf E}^{j_{m-1}}Y_{i_{m-1}},{\bf E}^{j_{m}}Y_{i_{m}})$
is an edge from the vertex
$(Y_{i_{0}},{\bf E}^{j_{1}}Y_{i_{1}},{\bf E}^{j_{2}}Y_{i_{2}},\ldots,{\bf E}^{j_{m-1}}Y_{i_{m-1}})$
to the vertex
$(Y_{i_{1}},{\bf E}^{j_{2}-j_{1}}Y_{i_{2}},\ldots,{\bf E}^{j_{m-1}-j_{1}}Y_{i_{m-1}},{\bf E}^{j_{m}-j_{1}}Y_{i_{m}}).$
The graph $G_{{\cal F},m}$ has $2^{nm-k}$ vertices and $2^{n(m+1)-k}$ edges. An Hamiltonian cycle in $G_{{\cal F},m}$ can be represented by a vertically periodic $2^{k}\times(2^{nm-k}+m-1)$ matrix in which each $n\times m$ matrix is contained exactly once. If the $2^{k}\times(m-1)$ matrix obtained from the first $m-1$ columns of this matrix equals (without a cyclic shift) to the $2^{k}\times(m-1)$ matrix obtained from the last $m-1$ columns, then the Hamiltonian cycle can be represented by a $2^{k}\times 2^{nm-k}$ doubly-periodic matrix. This condition is satisfied when $m>2$ or $n\neq k$ and the associated matrix forms a $(2^{k},2^{mn-k};n,m)$-DBAC with exactly one codeword, i.e., a $(2^{k},2^{mn-k};n,m)$-DBA. When $m=2$ and $n=k$ the last column will be shifted in $2^{n-1}$ positions compared to the first one and we cannot remove it to form a doubly-periodic array. This is something that has to be taken into account in our constructions too.
Similarly, an Eulerian cycle in $G_{{\cal F},m}$ can be represented by a vertically periodic $2^{k}\times(2^{n(m+1)-k}+m)$ matrix. If the $2^{k}\times m$ matrix obtained from the first $m$ columns of this matrix equals to the $2^{k}\times m$ matrix obtained from the last $m$ columns, then the cycle can be represented by a $2^{k}\times 2^{n(m+1)-k}$ doubly-periodic matrix. This is satisfied when $m>1$ or $n\neq k$ and the associated matrix forms a $(2^{k},2^{n(m+1)-k};n,m+1)$-DBA.
We are not interested in a code with one codeword as these codes are just de Bruijn arrays and it was proved by Paterson [42] that the necessary conditions of Theorem 1 are sufficient for this case. However, these codes can be used in our recursive constructions.
An $(2^{k},2^{t};n,m)$-DBAC, where the columns of the arrays are taken from a PF$(n,k)$, say ${\cal F}$, with $\Delta$ codewords can be generated from $G_{{\cal F},m}$. Such a code in which all the columns of the codewords are from ${\cal F}$ can be viewed as a factor in $G_{{\cal F},m}$. In this factor of $G_{{\cal F},m}$ the associated $2^{k}\times 2^{t}$ matrices must be doubly periodic. In other words, this factor in $G_{{\cal F},m}$ with $\Delta$ cycles of the same length $2^{t}$ forms a $(2^{k},2^{t};n,m)$-DBAC. Any cycle of length $2^{t}$ in $G_{{\cal F},m}$ can be represented by a $2^{k}\times(2^{t}+m-1)$ matrix. But, to be a codeword in a DBAC this is conditioned to the requirement that when a codeword is represented by its $2^{k}\times(2^{t}+m-1)$ matrix, the $2^{k}\times(m-1)$ matrix obtained from the first $m-1$ columns and the last $m-1$ columns of each matrix are equal. It should be noted that there are $(2^{k},2^{t};n,m)$-DBACs whose columns do not form a perfect factor in $G_{n}$. In all the direct constructions which will be presented in this paper the columns of the arrays will be cycles from a perfect factor. Moreover, the recursive constructions which be presented will make usage of the cycles in a perfect factor to determine when ${\bf D}_{0}^{-1}$ will be applied on columns of the array and when ${\bf D}_{1}^{-1}$ will be applied. This view of DBACs can be considered as a two-dimensional generalization for perfect factors. Moreover, the graph $G_{{\cal F},m}$ can be viewed as a two-dimensional generalization for the de Bruijn graph and a factor in the graph as a state diagram for a two-dimensional generalization for nonsingular feedback shift registers.
The generalization for two-dimensional feedback shift register using a perfect factor PF$(n,k)$ is straightforward. Assume that perfect factor ${\cal F}$ has $2^{n-k}$ cycles $X_{0},X_{1},\ldots,X_{2^{n-k}-1}$ of length $2^{k}$ and the register has $m$ cells $x_{1},x_{2},\ldots,x_{m}$. At each stage the cell $x_{\ell}$, $1\leq\ell\leq m$, stores a column vector $X_{i_{\ell}}$ of length $2^{k}$ which is a sequence of ${\cal F}$ in some shift ${\bf E}^{j_{\ell}}$ related to its zero state, where $j_{1}=0$, i.e., the sequence in $x_{1}$ is taken in its zero state. Using the feedback function $x_{m+1}=f(x)=f(x_{1},x_{2},\ldots,x_{m})$ we have that
$x_{m+1}=f(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{m}}X_{i_{m}}),$
i.e., $x_{m+1}$ equals to ${\bf E}^{j_{m+1}}X_{i_{m+1}}$, i.e., for each $\ell$, $11\leq\ell\leq m-$ 1 in the next stage controlled by a master clock $x_{\ell}$ receives its information from $x_{\ell+1}$ and hence it stores ${\bf E}^{j_{\ell+1}-j_{2}}X_{i_{\ell+1}}$. The shifts are taken modulo $2^{k}$, the length of the cycles in ${\cal F}$ and they are reduced by $j_{2}$ since in $x_{1}$ the sequence should be taken in its zero state. Finally $x_{m}$ will store ${\bf E}^{j_{m+1}-j_{2}}X_{i_{m+1}}$. The state diagram is a subgraph of $G_{{\cal F},m}$ and for the feedback function $f$ it has the edge
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{m}}X_{i_{m}})\longrightarrow(X_{i_{2}},{\bf E}^{j_{3}-j_{2}}X_{i_{3}},\ldots,{\bf E}^{j_{m}-j_{2}}X_{i_{m}},{\bf E}^{j_{m+1}-j_{2}}X_{i_{m+1}})~{}.$
Due to this definition and the state diagram of the feedback function we call the vertices of $G_{{\cal F},m}$ also states. As in the one-dimensional case we define a nonsingular two-dimensional feedback shift register to be a feedback shift register whose state diagram contains only cycles, i.e., it is a factor of $G_{{\cal F},m}$ associated with doubly-periodic matrices. We can now develop a theory for two-dimensional feedback shift registers in the same ways as was done in the one-dimensional case. Anyway, this complete theory is beyond the scope of the current paper, although some small fractions that are relevant to the current work will be mentioned.
The merge-or-split method is one feature that can be generalized from a factor in $G_{n}$ to a factor in $G_{{\cal F},m}$, i.e., from one-dimensional feedback shift register to two-dimensional feedback shift registers, where ${\cal F}$ is a factor in $G_{n}$ with cycles of length $2^{k}$. For simplicity we will consider factors in $G_{{\cal F},m}$ in which all the cycles have the same length $2^{t}$ and each associated $2^{k}\times 2^{t}$ matrix form an $(2^{k},2^{t};n,m)$-DBAC, i.e., the $2^{k}\times 2^{t}$ matrix is doubly periodic. A set of vertices in $G_{{\cal F},m}$ represented by $2^{k}\times m$ matrices is called a companion set if the $2^{k}\times(m-1)$ matrices defined by first $m-1$ columns of the matrices in the set are all equal. Two distinct vertices from this set will be called companion vertices (states) similarly to such pair of vertices in $G_{n}$. Two cycles which contain a pair of companion vertices from such a set, one vertex on each cycle, can be combined into one cycle when the merge-or-split method is applied on this pair of companion vertices in these two cycles. The associated $2^{k}\times 2^{t+1}$ matrix obtained from this join is doubly periodic since the sum of shifts in each cycle is 0 modulo $2^{k}$ which implies that sum of shifts in the combined cycle is also 0 modulo $2^{k}$. The reason is that the shifts in the two cycles are also the shifts of the combined cycle.
III Direct Constructions for de Bruijn Arrays Codes
In this section, we present five direct constructions for de Bruijn arrays codes. These constructions can be used for the recursive constructions presented in Section IV and for the merge-or-split method applied on some DBACs in Section V, to analyze the parameters of the $(2^{k},2^{t};n,m)$-DBACs that are obtained in our work. The first construction is aimed to construct $(2^{k},2^{t};n,1)$-DBACs for all possible parameters. The key idea is to notice that in each such array the columns are sequences of length $2^{k}$ such that each $n$-tuple is contained in exactly one of these columns. This implies that if the size of the code is $\Delta$, then the union of the sequences in all the columns of all the $\Delta$ arrays form a perfect factor in $G_{n}$ and vice versa. This perfect factor has $2^{n-k}$ cycles of length $2^{k}$, i.e., it is a PF$(n,k)$. Moreover, for this perfect factor $2^{t}$ can be any divisor of $2^{n-k}$. These observations imply the following theorem.
Theorem 5
An $(2^{k},2^{t};n,1)$-DBAC, $1\leq t\leq n-k$, of size $2^{n-k-t}$ exists if and only if there exists a PF$(n,k)$.
Corollary 2
An $(2^{k},2^{t};n,1)$-DBAC, $1\leq t\leq n-k$, of size $2^{n-k-t}$ exists if and only if $k\leq n<2^{k}$, i.e., $n<2^{k}\leq 2^{n}$.
Note, that when $t=0$, a $(2^{k},2^{t};n,1)$-DBAC is a perfect factor and when $t=n-k$, a $(2^{k},2^{t};n,1)$-DBAC is a de Bruijn array.
Next, we would like to have a direct construction for $(2^{k},2^{t};n,2)$-DBACs. For $t=1$ such codes do not exists by Lemma 1. Hence, we are intereted in the case where $t>1$. We notice that to construct such codes we can take a PF$(n,k)$, say ${\cal F}$, and guarantee that each two ordered sequences of ${\cal F}$ (not necessarily distinct) will be in all possible shifts one towards the other. This can be done in most cases as will be explained in the proof of the next theorem.
Theorem 6
If there exist a PF$(n,k)$, where $1\leq n-k$, then there exists a $(2^{k},2^{t};n,2)$-DBAC, where $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1 (if $t=2n-k$ then we have a $(2^{k},2^{t};n,2)$-DBA).
Proof:
Let ${\cal F}$ be a PF$(n,k)$ which by definition has $2^{n-k}$ cycles. To construct the appropriate code $\mathbb{C}$ of size $2^{2n-t-k}$ we have to show that we can arrange the sequences of ${\cal F}$ as columns in $2^{k}\times 2^{t}$ arrays such that each $n\times 2$ matrix is a contained as an $n\times 2$ window in exactly one of these arrays. Since each $n$-tuple appears as a window of length $n$ in one of the sequences of ${\cal F}$, it follows that each ordered two sequences of ${\cal F}$ (not necessarily distinct) must be in all the possible shifts between them. For each such a pair of sequences $({\cal S}_{1},{\cal S}_{2})$, we must have for each $i$, $10\leq i\leq 2^{k}-$ 1, the ordered pair $({\bf E}^{j}{\cal S}_{1},{\bf E}^{j+i}{\cal S}_{2})$ is contained as a pair of columns in exactly one of the codewords of $\mathbb{C}$ for some $j$. First, we will show that this is possible in the sense that the sum of all the shifts is 0 modulo $2^{k}$. Between any ordered pair of sequences the shifts are from 0 to $12^{k}-$ 1. The number of such ordered pairs is $2^{n-k}\cdot 2^{n-k}$. Hence, the total sum of all the shifts is
$\left(2^{n-k}\cdot 2^{n-k}\right)\sum_{i=0}^{2^{k}-1}i=2^{2(n-k)}(2^{k}-1)2^{k-1}\equiv 0~{}({\mbox{mod}}~{}2^{k})~{},$
since $n-k\geq 1$.
  1. 1.
    Assume first that $k=n-1$, i.e., ${\cal F}$ contains two sequences ${\cal A}$ and ${\cal B}$ of length $2^{n-1}$. We distinguish now between three cases depending on the value of $t$.
    Case 1: If $t=n$ (i.e., $t=k+1$), then there are exactly two codewords in $\mathbb{C}$ and they can be constructed as follows.
    $[{\cal A},{\bf E}{\cal B},{\cal A},{\bf E}^{2}{\cal B},{\cal A},{\bf E}^{3}{\cal B},\ldots,{\cal A},{\bf E}^{2^{t-1}-2}{\cal B},{\cal A},{\bf E}^{2^{t-1}-1}{\cal B},{\bf E}^{2^{t-1}-1}{\cal B},{\cal A}]$ (3)
    $[{\cal A},{\bf E}{\cal A},{\cal A},{\bf E}^{2}{\cal A},\ldots,{\cal A},{\bf E}^{2^{t-2}-1}{\cal A},{\cal A},{\bf E}^{2^{t-2}}{\cal A},{\bf E}^{2^{t-2}}{\cal B},{\cal B},{\bf E}^{2^{t-2}-1}{\cal B},{\cal B},\ldots,{\bf E}^{2}{\cal B},{\cal B},{\bf E}{\cal B},{\cal B}]~{}.$ (4)
    In the first codeword we only have the pairs $({\cal A},{\cal A})$ and $({\cal B},{\cal B})$ (we view the first sequence as being in its zero state) with the same two sequences in the pair. In the second codeword we have all the pairs of the form $({\cal A},{\bf E}^{i}{\cal A})$, where $1\leq i\leq 2^{t-2}$, and $({\bf E}^{i}{\cal A},{\cal A})$ which is equivalent to $({\cal A},{\bf E}^{-i}{\cal A})$, where $11\leq i\leq 2^{t-2}-$ 1. Since $k=t-1$ and the shifts are taken modulo $2^{k}$, it follows that each pair $({\cal A},{\bf E}^{i}{\cal A})$, where $0\leq i<2^{k}$, appears exactly once in one of the two codewords. Exactly the same arguments holds for the pairs of the form $({\cal B},{\bf E}^{i}{\cal B})$.
    In the second codeword we only have the pairs $({\cal A},{\cal B})$ and $({\cal B},{\cal A})$, where the two sequences in the pair are distinct. In the first codeword we have all the pairs of the form $({\cal A},{\bf E}^{i}{\cal B})$ and $({\bf E}^{i}{\cal B},{\cal A})=({\cal B},{\bf E}^{-i}{\cal A})$, where $11\leq i\leq 2^{t-1}-$ 1. Each such pair appears exactly once and since $k=t-1$, it follows that each required pair appears exactly once in one of the two codewords.
    Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
    Case 2: If $t=n-1$ (i.e., $t=k$), then there are exactly four codewords in $\mathbb{C}$. The first two codewords can be defined as in Eqs. (3) and (4), but the shifts are smaller since $t$ is smaller compared to $t$ of Case 1. This also implies that the number of columns in the codeword are smaller. The other two codewords can be defined as follows.
    $[{\cal B},{\bf E}^{2^{t-1}-1}{\cal A},{\cal B},{\bf E}^{2^{t-1}-2}{\cal A},{\cal B},{\bf E}^{2^{t-1}-3}{\cal A},\ldots,{\cal B},{\bf E}^{2}{\cal A},{\cal B},{\bf E}{\cal A},{\bf E}^{2^{k-1}+1}{\cal A},{\bf E}^{2^{k-1}}{\cal B}]$ (5)
    $[{\cal B},{\bf E}^{2^{t-1}-1}{\cal B},{\cal B},{\bf E}^{2^{t-1}-2}{\cal B},\ldots,{\cal B},{\bf E}^{2^{t-2}+1}{\cal B},{\cal B},{\bf E}^{2^{t-2}}{\cal B},{\bf E}^{2^{k-1}+2^{t-2}}{\cal A},{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-2}+1}{\cal A},$ (6)
    ${\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-2}+2}{\cal A},{\bf E}^{2^{k-1}}{\cal A},\ldots,{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-1}-2}{\cal A},,{\bf E}^{2^{k-1}}{\cal A},{\bf E}^{2^{k-1}+2^{t-1}-1}{\cal A},{\bf E}^{2^{k-1}}{\cal A}]~{}.$
    The same analysis as in Case 1 is done in this case, taken into account that $t=k$. We omit the specific discussions as a more comprehensive analysis will be made in Case 3, which is a generalization of this case. Hence, also in this case we have that $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
    Case 3: If $12\leq t<n-$ 1, then we have $2^{n-t+1}$ codewords. The first four codewords are defined exactly as in Eqs. (3), (4), (5), and (6).
    Note, that in these four codewords if a pair $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal A})$ was not used in one of these codewords, then also the pair $({\bf E}^{j+i}{\cal A},{\bf E}^{j}{\cal A})$ was not used in these codewords. The same it true for $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal B})$ and $({\bf E}^{j+i}{\cal B},{\bf E}^{j}{\cal B})$. As for the other pairs, if the pair $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal B})$ is contained in these codewords, then the pair $({\bf E}^{j+i}{\cal B},{\bf E}^{j}{\cal A})$ is also contained in these codewords.
    The analysis given for Case 1 also holds in the current case. Adding to it the pairs which are contained in the third and the fourth codewords defined in Eqs. (5) and (6), we have that the pairs which are contained in the four codewords are of the form $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal A})$, $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal B})$, $({\bf E}^{j}{\cal A},{\bf E}^{j+i}{\cal B})$, and $({\bf E}^{j}{\cal B},{\bf E}^{j+i}{\cal A})$, where $-2^{t-1}<i<2^{t-1}$ (in Case 2 we have that $t=k$ and hence also the pair with $i=-2^{t-1}=2^{t-2}$ is contained). Moreover, the pairs $({\bf E}^{j}{\cal A},{\bf E}^{j+2^{k-1}}{\cal A})$ and $({\bf E}^{j}{\cal B},{\bf E}^{j+2^{k-1}}{\cal B})$ are contained in the third codeword. The pairs $({\bf E}^{j}{\cal A},{\bf E}^{j+2^{k-1}}{\cal B})$ and $({\bf E}^{j}{\cal B},{\bf E}^{j+2^{k-1}}{\cal A})$ are contained in the fourth codeword.
    Now, it is straightforward to define the remaining codewords. There are $22^{n-t}-$ 2 codewords of the form
    $[{\cal A},{\bf E}^{i_{1}}{\cal B},{\cal A},{\bf E}^{i_{2}}{\cal B},{\cal A},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal B}],$ (7)
    where the $i_{j}$’s in all the $22^{n-t}-$ 2 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Finally, first there are $12^{n-t-1}-$ 1 codewords of the form
    $[{\cal A},{\bf E}^{i_{1}}{\cal A},{\cal A},{\bf E}^{i_{2}}{\cal A},{\cal A},{\bf E}^{i_{3}}{\cal A},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal A}]$ (8)
    where the $i_{j}$’s in all the $12^{n-t-1}-$ 1 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Secondly, there are also $12^{n-t-1}-$ 1 codewords of the form
    $[{\cal B},{\bf E}^{i_{1}}{\cal B},{\cal B},{\bf E}^{i_{2}}{\cal B},{\cal B},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal B},{\bf E}^{i_{2^{t-1}}}{\cal B}]$ (9)
    where the $i_{j}$’s in all the $12^{n-t-1}-$ 1 codewords are all the distinct integers between $2^{t-1}$ and $2^{k}-2^{t-1}$ in any order, except for $2^{k-1}$.
    Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
  2. 2.
    For $k<n-1$, the PF$(n,k)$ contains $2^{n-k}\geq 4$ sequences, ${\cal A}_{1},{\cal A}_{2},\ldots,{\cal A}_{2^{n-k}}$. These sequences are ordered in pairs $\{{\cal A}_{i},{\cal A}_{i+1}\}$, where $i$ is an odd integer between 1 and $12^{n-k}-$ 1. For each pair we construct in $\mathbb{C}$ the same codewords as was done in Eqs. (3)-(9) for ${\cal A}$ and ${\cal B}$. The analysis of the obtained pairs is done exactly as was done in Case 3.
    It remains to form the pairs between ${\cal A}_{i}$ and ${\cal A}_{j}$, where $i<j$ and $\{{\cal A}_{i},{\cal A}_{j}\}\neq\{A_{2\ell+1},A_{2\ell+2}\}$, $10\leq\ell\leq 2^{n-k-1}-$ 1. For each such pair $\{{\cal A}_{i},{\cal A}_{j}\}$ (for simplicity this pair is denoted by $\{{\cal A},{\cal B}\}$), we form the following $2^{k-t+1}$ codewords (when $k\geq t$).
    $[{\cal A},{\bf E}^{i_{1}}{\cal B},{\cal A},{\bf E}^{i_{2}}{\cal B},{\cal A},{\bf E}^{i_{3}}{\cal B},\ldots,{\cal A},{\bf E}^{i_{2^{t-1}}}{\cal B}],$
    where the $i_{j}$’s in all the $2^{k-t+1}$ codewords are all the distinct integers between 00 and $12^{k}-$ 1 in any order.
    This implies that $\mathbb{C}$ is a $(2^{k},2^{t};n,2)$-DBAC.
Thus, for each set of parameters a $(2^{k},2^{t};n,2)$-DBAC, where $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1, was constructed. ∎
Corollary 3
If $k<n<2^{k}$ and $12\leq t\leq 2n-k-$ 1, then there exists a $(2^{k},2^{t};n,2)$-DBAC.
Theorem 6 does not cover the case where $k=n$, i.e., when the perfect factor contains only a span $n$ de Bruijn sequence. The technique that was used will not work in this case since $\sum_{i=0}^{2^{n}-1}i\equiv 2^{n-1}~{}({\mbox{mod}}~{}2^{n})$ and hence we can have that in at least one of the constructed arrays, the sum of all the shifts will be different from 0 modulo $2^{n}$ as required.
The next theorem and its construction is a generalization of a technique presented by Ma [31], where in [31] one sequence over $\mathbb{Z}_{2^{m}}$ is used, while we use all the sequences of a perfect factor over $\mathbb{Z}_{2^{m}}$.
Theorem 7
[…]
Proof:
[…]
$[{\cal S},{\bf E}^{q_{0}}{\cal S},{\bf E}^{q_{0}+q_{1}}{\cal S},{\bf E}^{q_{0}+q_{1}+q_{2}}{\cal S},\ldots,{\bf E}^{\sum_{i=0}^{2^{t}-2}q_{i}}{\cal S}].$
The generated codeword form a $2^{n}\times 2^{t}$ matrix and since $\sum_{i=0}^{2^{t}-1}q_{i}\equiv 0~{}({\mbox{mod}}~{}2^{n})$, it follows that the matrix is doubly periodic. The perfect factor ${\cal F}$ has $2^{nm-t}$ sequences of length $2^{t}$ and each $m$-tuple is contained as a window of length $m$ in exactly one of the sequences of ${\cal F}$. From each sequence of ${\cal F}$ we generate one codeword in $\mathbb{C}$ and hence the total number of windows in the codewords of $\mathbb{C}$ is $2^{nm-t}\cdot 2^{n}\cdot 2^{t}=2^{n(m+1)}$. Therefore, to prove that the generated code is a $(2^{n},2^{t};n,m+1)$-DBAC it is sufficient to show that each $n\times(m+1)$ matrix is a window in one of the codewords of $\mathbb{C}$. Let $A$ be such an $n\times(m+1)$ matrix. Each column of $A$ is an $n$-tuple in ${\cal S}$. Hence, we have that $A$ is contained in the $2^{n}\times(m+1)$ vertically periodic array
${\cal S},{\bf E}^{b_{1}}{\cal S},{\bf E}^{b_{2}}{\cal S},\ldots,{\bf E}^{b_{m}}{\cal S},$
for some $m$-tuple $(b_{1},b_{2},\ldots,b_{m})$ over $\mathbb{Z}_{2^{n}}$. Consider now another $m$-tuple $\beta$ with entries from $\mathbb{Z}_{2^{n}}$, where
$\beta=(b_{1},b_{2}-b_{1},b_{3}-b_{2},b_{4}-b_{3},\ldots,b_{m}-b_{m-1})=(p_{j},p_{j+1},\ldots,p_{j+m-1}).$
Since each $m$-tuple is contained in a window of length $m$ in one sequence of ${\cal F}$, it follows that $\beta$ is contained as such a window in such a sequence. The associated codeword in $\mathbb{C}$ has the following $m+1$ successive columns
${\bf E}^{\ell}{\cal S},{\bf E}^{\ell+p_{j}}{\cal S},{\bf E}^{\ell+p_{j}+p_{j+1}}{\cal S},\ldots,{\bf E}^{\ell+\sum_{i=0}^{m-1}p_{j+i}}{\cal S}={\bf E}^{\ell}{\cal S},{\bf E}^{\ell+b_{1}}{\cal S},{\bf E}^{\ell+b_{2}}{\cal S},\ldots,{\bf E}^{\ell+b_{m}}{\cal S}~{},$
for some $\ell$. These $m+1$ consecutive columns contain the matrix $A$ in $n$ consecutive rows. Thus, $\mathbb{C}$ is a $(2^{n},2^{t};n,m+1)$-DBAC. ∎
[…]
Theorem 8
Let ${\cal F}$ be a PF$(n,k)$, $m=2^{t}-1$, $t\geq k$, and define the following set of matrices
$\mathbb{C}\triangleq\{[X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{m}}X_{i_{m}},{\bf E}^{j_{m+1}}X_{i_{m+1}}]~{}:~{}$
$X_{i_{r}}\in{\cal F},~{}1\leq r\leq m+1,~{}\sum_{r=1}^{m+1}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k}),~{}\sum_{r=2}^{m+1}j_{r}\equiv 0~{}({\mbox{mod}}~{}2^{k}),~{}$
$0\leq j_{r}\leq 2^{k}-1,~{}2\leq r\leq m+1\},$
where each $X_{i_{r}}$ is a column vector and its cyclic shifts is related to its zero state. Then, the code $\mathbb{C}$ is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$.
Proof:
The codewords of the construction of $\mathbb{C}$ are $2^{k}\times 2^{t}$ matrices, but they can be also viewed as cycles of length $2^{t}$ in the graph $G_{{\cal F},m}$. We start by showing that the horizontal period of each cycle defined in $\mathbb{C}$ is ${m+1=2^{t}}$, i.e., the associated matrix does not have a smaller horizontal periodicity. If the period of some cycle ${\cal C}$ in $\mathbb{C}$ is smaller than $2^{t}$, i.e., the associated matrix has a smaller horizontal periodicity, then it should be a divisor of $2^{t}$, i.e., a power of 2. Assume for the contrary that such a matrix has a horizontal periodicity, i.e., it has period $2^{s}$ for some $s$, ${0\leq s<t}$ which implies that it consists of $2^{s}$ disjoint $2^{k}\times 2^{t-s}$ equal vertically periodic matrices. Since $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$, it follows that $\sum_{r=1}^{2^{t}}i_{r}=\alpha 2^{n-k}+1$, for some positive integer $\alpha$, and hence $\sum_{r=1}^{2^{s}}i_{r}=\frac{\alpha 2^{n-k}+1}{2^{t-s}}$. But, $\frac{\alpha 2^{n-k}+1}{2^{t-s}}$ is not an integer, a contradiction. Therefore, all the cycles of $\mathbb{C}$ have no smaller horizontal periodicity and hence they form $2^{k}\times 2^{t}$ arrays with period $2^{t}$.
We continue by computing the size of the code $\mathbb{C}$. Each $X_{i_{r}}$, $1\leq r\leq m$, is taken arbitrarily from ${\cal F}$ and hence it can be chosen in $2^{n-k}$ distinct ways. The sequence $X_{i_{m+1}}$ is determined by the equation ${\sum_{r=1}^{m+1}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})}$. The first cycle $X_{i_{1}}$ is taken in its zero state, while the next $m-1$ cycles can be taken in any of their $2^{k}$ possible cyclic shifts. The shift of the last cycle $X_{i_{m+1}}$ is determined by the equation $\sum_{r=2}^{m+1}j_{r}\equiv 0~{}({\mbox{mod}}~{}2^{k})$. Hence, each one of the middle $m-1$ cycles can be chosen in $2^{n-k}2^{k}=2^{n}$ distinct ways. The first cycle can be chosen in $2^{n-k}$ distinct ways and the last cycle and its shift are determined by the first $m$ cycles and their shifts. Therefore, there are together $2^{n(m-1)}2^{n-k}=2^{nm-k}$ distinct ways to choose all the cycles and their shifts.
Now, we calculate the number of cycles which are counted more than once in this enumeration. Each constructed cycle ${\cal C}$ can start with any one of its $m+1=2^{t}$ columns. Such a column, say the $r$-th column, will be taken in the shift where its zero state is the first $n$-tuple. The other columns are defined by the order of the columns in ${\cal C}$ and their shifts are taken exactly as in ${\cal C}$ related to the first state of the new first column. Therefore, each codeword ${\cal C}\in\mathbb{C}$ is constructed exactly $2^{t}$ times in the construction of $\mathbb{C}$. Thus, the size of $\mathbb{C}$ is $2^{nm-k}/2^{t}=2^{nm-k-t}$. Note, that the sum of all the shifts, in this new arrangement of the first column, will remain congruent to 0 modulo $2^{k}$ since $t\geq k$ and hence the sum of the new shifts will be changed by $j_{r}2^{t}$ modulo $2^{k}$ which is congruent to 0 modulo $2^{k}$, where the $r$-th column of ${\cal C}$ is the new first column.
Therefore, the number of $n\times m$ windows in the all the codewords of $\mathbb{C}$ is $2^{nm-k-t}\cdot 2^{k}\cdot 2^{t}=2^{nm}$. Hence, to complete the proof it is sufficient to show that each $n\times m$ matrix appears as a window in a codeword of $\mathbb{C}$. Consider such an $n\times m$ matrix $A$. The columns of $A$ are $n$-tuples in cycles of ${\cal F}$ and hence these cycles can be arranged in a codeword ${\cal C}\in\mathbb{C}$ with all the necessary shifts as required by $A$. This arrangements can be guaranteed by the definition of the first $m=2^{t}-1$ columns in a codeword of $\mathbb{C}$ to form a window with the matrix $A$ in ${\cal C}$, since there are no constraints on the shifts in the construction for the $m-1$ columns which follows the first one.
Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$. ∎
[…]
Theorem 9
Let ${\cal F}$ be a PF$(n,k)$, where ${\cal F}$ does not contain self-dual sequences and, ${\cal S}\in{\cal F}$ if and only if $\bar{{\cal S}}\in{\cal F}$. Define the following set of matrices
$\mathbb{C}\triangleq\{[X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{\ell}}X_{i_{\ell}},\bar{X}_{i_{1}},{\bf E}^{j_{2}}\bar{X}_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}\bar{X}_{i_{2^{t}}}]~{}:~{}$
$X_{i_{r}}\in{\cal F},~{}1\leq r\leq 2^{t},~{}0\leq j_{r}\leq 2^{k}-1,~{}2\leq r\leq 2^{t}\},$
where each $X_{i_{1}}$, $1\leq r\leq 2^{t}$, is a column vector and its cyclic shift is related to its zero state. Then, the code $\mathbb{C}$ is a $(2^{k},2^{t+1};n,2^{t})$-DBAC of size $2^{n2^{t}-k-t-1}$.
Proof:
The codewords of the construction are $2^{k}\times 2^{t+1}$ doubly-periodic matrices, but they can also be viewed as cycles of length $2^{t+1}$, in $G_{{\cal F},2^{t}}$. We start by showing that horizonal period of each codeword defined in $\mathbb{C}$ is $2^{t+1}$. If the period of some codeword ${\cal C}$ in $\mathbb{C}$ is smaller than $2^{t+1}$, then it should be a divisor of $2^{t}$, i.e., a power of 2. but this will imply that the sequence $X_{i_{1}}$ is some shift of the sequence $\bar{X}_{i_{1}}$, i.e., this sequence is self-dual which is impossible by our choice of ${\cal F}$. Therefore, all the codewords of $\mathbb{C}$ have no smaller horizontal periodicity and hence they form $2^{k}\times 2^{t+1}$ doubly-periodic arrays with no smaller horizontal periodicity.
We continue by computing the size of the code $\mathbb{C}$. Each $X_{i_{r}}$, $1\leq r\leq 2^{t}$, is taken arbitrarily from ${\cal F}$ and hence it can be chosen in $2^{n-k}$ distinct ways. The first cycle $X_{i_{1}}$ is taken in its zero state, while the next $12^{t}-$ 1 cycles can be taken in any of their $2^{k}$ cyclic shifts. The last $2^{t}$ cycles and their shifts are determined by the first $2^{t}$ cycles, i.e., these are complement cycles with the same shifts. Hence, each of the first $2^{t}$ cycles, except for $X_{i_{1}}$, has $2^{n-k}2^{k}=2^{n}$ possible distinct choices. Therefore, there are $2^{n(2^{t}-1)}\cdot 2^{n-k}=2^{n2^{t}-k}$ distinct choices for all the codewords in $\mathbb{C}$ and their shifts.
Now, we calculate the number of codewords of $\mathbb{C}$ which are counted more than once in this enumeration. Each constructed cycle ${\cal C}$ can start with any one of its $2^{t+1}$ columns. This column will be taken in the shift where its zero state is the first $n$-tuple. The other columns are defined by the order of the columns in ${\cal C}$ and their shifts are taken exactly as in ${\cal C}$ related to the first state of the new first column. Therefore, each codeword ${\cal C}$ is constructed exactly $2^{t+1}$ times in the construction of $\mathbb{C}$. Thus, the size of $\mathbb{C}$ is $2^{n2^{t}-k}/2^{t+1}=2^{n2^{t}-k-t-1}$.
Therefore, the number of $n\times 2^{t}$ windows in the all the codewords of $\mathbb{C}$ is $2^{n2^{t}-k-t-1}\cdot 2^{k}\cdot 2^{t+1}=2^{n2^{t}}$. Hence, to complete the proof it is sufficient to show that each $n\times 2^{t}$ matrix appears as a window in a codeword of $\mathbb{C}$. Consider such a matrix $A$. The columns of $A$ are $n$-tuples in cycles of ${\cal F}$ and hence these cycles can be rearranged in a codeword ${\cal C}$ to form the window with the matrix $A$ since there are no constraints in the construction on the order of $2^{t}$ consecutive columns.
Thus, $\mathbb{C}$ is a $(2^{k},2^{t+1};n,2^{t})$-BBAC of size $2^{n2^{t}-k-t-1}$. ∎
[…]
The following result which is a consequence from the discussion in Section II-D implies that the requirement for a perfect factor in Theorem 9 can be always satisfied.
Lemma 12
[…]
Proof:
The methods used to generate perfect factors in Theorem 4 holds also to obtain perfect factors without self-dual sequences and if ${\cal F}$ is the PF$(n,k)$, then ${\cal S}\in{\cal F}$ if and only if $\bar{{\cal S}}\in{\cal F}$. The only exception is when $n=2^{k-1}$ and the sequences which are obtained in PF$(n,k)$ are with linear complexity $12^{k-1}+$ 1 and hence by Lemma 7 the obtained sequences are self-dual.
[…]
IV Recursive Constructions for de Bruijn Arrays Codes
For a recursive construction, we are given a $(2^{k},2^{t};n,m)$-DBAC $\mathbb{C}$ with some parameters which will be specified in the construction and a perfect factor ${\cal F}$ with certain parameters which depend on the parameters of $\mathbb{C}$. The sequences of ${\cal F}$ are used as indicators associated with the columns of the codewords of $\mathbb{C}$, from which we decide whether to apply ${\bf D}^{-1}_{0}$ or to apply ${\bf D}^{-1}_{1}$. All the recursive constructions which will be presented are generalizations of two constructions presented by Fan, Fan, Ma, and Siu [17].
Theorem 10
If there exists a $(2^{k},2^{m};n,m)$-DBAC whose size is $\Delta$, where $m>1$, and its columns are sequences of period $2^{k}$ and even weight, then there exists an $(2^{k},2^{m};n+1,m)$-DBAC of size $2^{m}\Delta$.
Proof:
Let $\mathbb{C}$ be a $(2^{k},2^{m};n,m)$-DBAC of size $\Delta$ whose columns are sequences of period $2^{k}$ and even weight. Let $A=[{\cal A}_{1}{\cal A}_{2}~{}\cdots~{}{\cal A}_{2^{m}}]$ be any codeword (a $2^{k}\times 2^{m}$ doubly-periodic array) of $\mathbb{C}$. By the discussion in Section II-C we infer that by applying ${\bf D}^{-1}$ on a column from $A$, the outcome is two complementary sequences of the same period $2^{k}$. These sequences will be used to construct a new $2^{k}\times 2^{m}$ doubly-periodic array $B$ which will be part of a new code $\mathbb{C}^{\prime}$. Let
${\cal S}=[s_{1},s_{2},\ldots,s_{2^{m}}]$
be a span $m$ de Bruijn sequence. For the code $\mathbb{C}^{\prime}$ we define a new array $B$. Let the $i$-th column of $B$ be defined by, $B_{i}={\bf D}^{-1}_{s_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of $A$ and $1\leq i\leq 2^{m}$. Clearly, $B$ is an $2^{k}\times 2^{m}$ doubly-periodic array. The same method is applied with the $2^{m}$ distinct cyclic shifts of ${\cal S}$. Since the code $\mathbb{C}$ is of size $\Delta$ and each array $A$ is used to construct $2^{m}$ new arrays, it follows that the new code $\mathbb{C}^{\prime}$ has $2^{m}\Delta$ codewords. Each generated codeword in $\mathbb{C}^{\prime}$ has $2^{k+m}$ distinct windows of size $(n+1)\times m$ for a total of $2^{k+m}2^{m}\Delta$ windows of size $(n+1)\times m$ in all the arrays of $\mathbb{C}$. Since $2^{k+m}2^{m}\Delta=2^{nm}2^{m}=2^{(n+1)m}$ (note that $2^{k+m}\Delta=2^{nm}$ because $\mathbb{C}$ is a $(2^{k},2^{m};n,m)$-DBAC with $\Delta$ codewords), it follows that to complete the proof it is sufficient to show that all the $(n+1)\times m$ windows in the codewords of $\mathbb{C}^{\prime}$ are distinct.
Assume for the contrary that there exist two $(n+1)\times m$ windows $X$ and $Y$ in two codewords of $\mathbb{C}^{\prime}$, $B_{1}$ and $B_{2}$ (not necessarily distinct) obtained from two codewords $A_{1}$ and $A_{2}$ of $\mathbb{C}$ (not necessarily distinct), respectively, for which $X=Y$ (if $A_{1}=A_{2}$ then $X$ and $Y$ are windows in distinct positions). Let ${\cal X}$ and ${\cal Y}$ be the two $n\times m$ arrays obtained from $X$ and $Y$, respectively, by applying ${\bf D}$ on the columns of $X$ and $Y$, respectively (note that these columns are acyclic sequences). Clearly, if $X=Y$, then ${\cal X}={\cal Y}$. Since $A_{1}$ and $A_{2}$ are codewords in the $(2^{k},2^{m};n,m)$-DBAC $\mathbb{C}$, it follows that all the $n\times m$ windows in all its codewords are distinct and hence ${\cal X}={\cal Y}$ implies that they are obtained from the same position of a codeword $A$, i.e., $A_{1}=A_{2}$, using different shifts of ${\cal S}$. Moreover, ${\cal X}={\cal Y}$ implies that they are in columns of $B_{1}$ and $B_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple in ${\cal S}$. But, this implies that $B_{1}=B_{2}$ and ${\cal S}$ is taken with the same shift to obtain $A$ and hence $X$ and $Y$ are the same $(n+1)\times m$ window with the same position in $A$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k},2^{m};n+1,m)$-DBAC of size $2^{m}\Delta$. ∎
What DBAC with new parameters can be obtained using Theorem 10? The direct constructions were used to form $(2^{k},2^{t};n,m)$-DBACs, where the columns of the codewords are sequences from an PF$(n,k)$ say ${\cal F}$. The construction of Theorem 10 implies the existence of a $(2^{k},2^{t};n+1,m)$-DBAC. But, such a DBAC can be obtained using a direct construction with a PF$(n+1,k)$ that can be obtained from ${\bf D}^{-1}{\cal F}$. This implies that with this recursion we cannot obtain DBACs with new parameters. On the other hand we can transpose the obtained DBACs by the direct constructions and use the recursion to obtained DBACs with new parameters. For this purpose we have to examine the weight of the rows in the direct constructions. To demonstrate the analysis we concentrate on two of the constructions, those obtained from Theorems 8 and 9.
For the construction in Theorem 8, we can have both rows of even weight and odd weight in the construction and hence the transpose will not be of much help. For the construction in Theorem 9, the codewords are $2^{k}\times 2^{t}$ matrices, where the $i$-th column is the complement of the $(2^{t-1}+i)$-th column. In other words, each row is a self-dual sequence whose linear complexity is $12^{t-1}+$ 1 which implies that its weight is even. Moreover, applying ${\bf D}^{-\ell}$ on the sequences up to $\ell\leq 2^{t-1}-2$ times yield sequences of even weight, so definitely we will obtain DBCAs with new parameters with the construction of Theorem 9. Another iteration of ${\bf D}^{-1}$ yields sequences with linear complexity $2^{t}$ which are sequences of odd weight. This makes this construction very important in the other recursive constructions.
Theorem 11
If there exists a $(2^{k},2^{m-\ell};n,m)$-DBAC, where $\ell\geq 0$, whose size is $\Delta$, where $m<2^{m-\ell}$, and the columns of the codewords are sequences of period $2^{k}$ and even weight, then there exists an $(2^{k},2^{m-\ell};n+1,m)$-DBAC of size $2^{m}\Delta$.
Proof:
The construction of a code $\mathbb{C}^{\prime}$ from a $(2^{k},2^{m-\ell};n,m)$-DBAC $\mathbb{C}$ and the correctness of the construction is almost the same as was done in the proof of Theorem 10. The main distinction is that in Theorem 10 one span $m$ de Bruijn sequence ${\cal S}$ is used for the construction. Instead of ${\cal S}$, we take the span $m$ sequences of ${\cal F}$, a PF$(m,m-\ell)$, which exists by Theorem 2. We apply the same technique as was applied in Theorem 10 with ${\cal S}$, this time with all the sequences of ${\cal F}$, each one is taken as ${\cal S}$ in all its cyclic shifts. ∎
The construction of Theorem 11 is a generalization of the one in Theorem 10 and the analysis for new parameters should be the same.
Theorem 12
If there exists a $(2^{k},2^{t};n,m)$-DBAC whose size is $\Delta$ and its columns are sequences of period $2^{k}$ and even weight, then there exists an $(2^{k},2^{m+t};n+1,m)$-DBAC of size $\Delta$.
Proof:
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC and let $A$ be a codeword in $\mathbb{C}$ whose columns are sequences with period $2^{k}$ and even weight. Define
${\cal A}\triangleq[\overbrace{A~{}A~{}\cdots~{}A}^{2^{m}~{}\text{copies~{}of}~{}A}].$
Clearly, ${\cal A}$ is an $2^{k}\times 2^{m+t}$ array. If we apply the operator ${\bf D}^{-1}$ on a column from $A$, the outcome is two complementary sequences of the same length $2^{k}$. These sequences will be used to construct a new array ${\cal B}$ in a new code $\mathbb{C}^{\prime}$. Let
${\cal S}=[s_{0},s_{1},\ldots,s_{2^{m}-1}]$
be a span $m$ de Bruijn sequence whose first $m$ bits are zeros and hence the sequence $[s_{1},s_{2},\ldots,s_{2^{m}-1}]$ is the shortened de Bruijn sequence obtained from ${\cal S}$ (starts with $m-1$ zeros). Define the sequence
$B\triangleq(\overbrace{0,0,\ldots,0}^{2^{t}~{}\text{zeroes}},\overbrace{s_{1},s_{2},\ldots,s_{2^{m}-1},\ldots,s_{1},s_{2},\ldots,s_{2^{m}-1}}^{2^{t}~{}\text{copies~{}of}~{}s_{1},s_{2},\ldots,s_{2^{m}-1}})$
of length $2^{t}+(2^{m}-1)2^{t}=2^{m+t}$ and denote $B=(b_{1},b_{2},\ldots,b_{2^{m+t}})$.
We define a codeword ${\cal B}$ in a new code $\mathbb{C}^{\prime}$. Let the $i$-th column of ${\cal B}$ be defined by, ${\cal B}_{i}={\bf D}^{-1}_{b_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of ${\cal A}$ and $1\leq i\leq 2^{m+t}$. The codeword ${\cal B}$ is an $2^{k}\times(2^{m+t})$ doubly-periodic array and hence it has $2^{k+m+t}$ windows of size $(n+1)\times m$. Since $2^{k+m}2^{t}\Delta=2^{m}2^{nm}=2^{(n+1)m}$ (since $2^{k}2^{t}\Delta=2^{nm}$), it follows that to complete the proof it is sufficient to show that each $(n+1)\times m$ binary array is contained at most once as an $(n+1)\times m$ window in some ${\cal B}$.
Assume for the contrary that there exist two $(n+1)\times m$ windows ${\cal X}$ and ${\cal Y}$ in two codewords ${\cal B}_{1}$ and ${\cal B}_{2}$ of $\mathbb{C}^{\prime}$ (not necessary distinct) for which ${\cal X}={\cal Y}$. Let $X$ and $Y$ be the two $n\times m$ arrays obtained from ${\cal X}$ and ${\cal Y}$, respectively, by applying ${\bf D}$ on the columns of ${\cal X}$ and ${\cal Y}$, respectively (note that these columns are acyclic sequences). Clearly, if ${\cal X}={\cal Y}$, then $X=Y$. Since $A$ is a codeword in an $(2^{k},2^{t};n,m)$-DBAC, it follows that all its $n\times m$ windows are distinct and hence $X=Y$ implies that they are obtained from the same position of $A$ in different locations of ${\cal A}$. Moreover, $X=Y$ implies that ${\cal X}$ and ${\cal Y}$ are in columns of ${\cal B}_{1}$ and ${\cal B}_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple in $B_{1}$ and $B_{2}$. The fact that g.c.d.$(2^{t},2^{m}-1)=1$ implies that the projection of two sets of $m$ columns on ${\cal A}$ and the associated two sets of $m$ entries in $B_{1}$ and $B_{2}$ (where the two sets of columns from ${\cal A}$ yield the same $2^{k}\times m$ window) cannot yield the same pair of $m$-tuples from $B_{1}$ and $B_{2}$ and hence cannot yield the same pair of $(n+1)\times m$ sub-matrices in ${\cal B}_{1}$ and ${\cal B}_{2}$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is a $(2^{k},2^{m+t};n+1,m)$-DBAC of size $\Delta$. ∎
The recursive construction in Theorem 12 yield DBCAs with new parameters as the horizontal length of an array is getting much larger without enlarging the horizontal length of the windows. Moreover, it can be also works on the transposed arrays as the constructions in the previous two theorems.
Theorem 13
If there exists a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$ whose columns are sequences of period $2^{k}$ and even weight, then there exists a $(2^{k},2^{m+t-1};n+1,m)$-DBAC of size $2\Delta$.
Proof:
The proof is essentially the same as the one of Theorem 12 with one difference. In Theorem 12 a span $m$ de Bruijn sequence is used. In the current theorem we use a span $m$ half de Bruijn sequence ${\cal S}$ which starts with $m$ zeros. In Theorem 12 $A$ is taken $2^{m}$ times to form ${\cal A}$. In the proof of the current theorem, $A$ is taken $2^{m-1}$ times to form ${\cal A}$. Moreover, we use the complement of ${\cal S}$, $\bar{{\cal S}}$ instead of ${\cal S}$ with ones instead of zeros to form another set of codewords. Similarly, instead of a sequence $B$ we use two sequences, $B$ and $\bar{B}$. ∎
The analysis for the recursive construction of Theorem 13 is very similar to the one of Theorem 12.
Theorem 14
If there exists a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where $3\leq m<2^{t}$ and $2\leq n<2^{k}$, whose columns are sequences of odd weight, then there exists a $(2^{k+1},2^{t};n+1,m)$-DBAC of size $2^{m-1}\Delta$.
Proof:
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$ whose columns are sequences of period $2^{k}$ and odd weight. Let $A=[{\cal A}_{1}{\cal A}_{2}~{}\cdots~{}{\cal A}_{2^{t}}]$ be any codeword (a $2^{k}\times 2^{t}$ doubly-periodic array) of $\mathbb{C}$. By the discussion in Section II-C, applying ${\bf D}^{-1}$ on a column from $A$, the outcome is a self-dual sequence of length $2^{k+1}$. These sequences will be used to construct a new array $B$ which will be part of $\mathbb{C}^{\prime}$. Let ${\cal F}$ be a PF$(m,t)$ with no self-dual sequences, where ${\cal C}\in{\cal F}$ if and only if $\bar{{\cal C}}\notin{\cal F}$. For each such pair of sequences $\{{\cal C},\bar{{\cal C}})$, let ${\cal S}={\cal C}$ or ${\cal S}=\bar{{\cal C}}$ and let
${\cal S}=[s_{1},s_{2},\ldots,s_{2^{t}}].$
For the code $\mathbb{C}^{\prime}$ we define a new doubly-periodic array $B$. Let the $i$-th column of $B$ be defined by, $B_{i}={\bf D}^{-1}_{s_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of $A$ and $1\leq i\leq 2^{m-1}$. Clearly, $B$ is a $(2^{k+1})\times 2^{t}$ array. We apply the same method with the $2^{t}$ distinct cyclic shifts of ${\cal S}$. Since $\mathbb{C}$ is of size $\Delta$ and each array $A$ is used to construct $2^{m-t-1}2^{t}=2^{m-1}$ new arrays in $\mathbb{C}^{\prime}$, it follows that $\mathbb{C}^{\prime}$ has $2^{m-1}\Delta$ codewords. Each codeword of $\mathbb{C}^{\prime}$ has $2^{k+t+1}$ distinct windows of size $(n+1)\times m$ for a total of $2^{m-1}\Delta 2^{k+t+1}$ windows of size $(n+1)\times m$ in all the arrays. Since $2^{m-1}\Delta 2^{k+t+1}=2^{nm}2^{m}=2^{(n+1)m}$ (because $\Delta 2^{k+t}=2^{nm}$), it follows that to complete the proof it is sufficient to show that each $(n+1)\times m$ binary array is contained at most once as an $(n+1)\times m$ window in one codeword of $\mathbb{C}^{\prime}$.
Assume for the contrary that there exist two $(n+1)\times m$ windows $X$ and $Y$ in two codewords of $\mathbb{C}^{\prime}$, $B_{1}$ and $B_{2}$ (not necessarily distinct) obtained from two codewords $A_{1}$ and $A_{2}$ of $\mathbb{C}$ (not necessarily distinct), respectively, for which $X=Y$. Let ${\cal X}$ and ${\cal Y}$ be the two $n\times m$ arrays obtained from $X$ and $Y$, respectively, by applying ${\bf D}$ on the columns of $X$ and $Y$, respectively (note that these columns are acyclic sequences). Clearly, if $X=Y$, then ${\cal X}={\cal Y}$. Since $A_{1}$ and $A_{2}$ are codewords in an $(2^{k},2^{t};n,m)$-DBAC, it follows that all the $n\times m$ windows in all its codewords are distinct and hence ${\cal X}={\cal Y}$ implies that they are obtained from the same position of a codeword $A$, i.e., $A_{1}=A_{2}$, using shifts of two sequences ${\cal S}_{1}$ and ${\cal S}_{2}$. Moreover, ${\cal X}={\cal Y}$ implies that they are in columns of $B_{1}$ and $B_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple ${\cal S}$ (either ${\cal S}_{1}$ or ${\cal S}_{2}$). But, this implies that $B_{1}=B_{2}$ and ${\cal S}$ is taken with the same shift to obtain $A$ and hence $X$ and $Y$ are the same window with the same position in $A$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k+1},2^{t};n+1,m)$-DBAC of size $2^{m-1}\Delta$. ∎
What DBACs with new parameters can be obtained using Theorem 14? The analysis is very similar to the one in Theorem 10. Applying the recursive construction on the DBACs obtained from the direct constructions will not yield DBACs with new parameters. Applying it on the transpose will also not help since the rows of the arrays either have even weight or have some rows of odd weight and some rows of even weight. The construction can be used to obtain DBACs with new parameters if it will work on DBACs obtained from the previous recursive constructions once the columns will be of odd weight.
Theorem 15
If there exists a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$ whose columns are sequences of odd weight, then there exists a $(2^{k+1},2^{m+t-1};n+1,m)$-DBAC of size $\Delta$.
Proof:
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC and $A$ a codeword in $\mathbb{C}$ whose columns are sequences with period $2^{k}$ and odd weight and define
${\cal A}\triangleq[\overbrace{A~{}A~{}\cdots~{}A}^{2^{m-1}~{}\text{copies~{}of}~{}A}].$
The codeword ${\cal A}$ is an $2^{k}\times 2^{m+t-1}$ doubly-periodic array. If we apply the operator ${\bf D}^{-1}$ on a column from $A$, the outcome is a self-dual sequence length $2^{k+1}$. These sequences will be used to construct a new array ${\cal B}$, where all these sequences have period $2^{k+1}$. Let
${\cal S}=[s_{0},s_{1},\ldots,s_{2^{m-1}-1}]$
be a span $m$ half de Bruijn sequence whose first $m$ bits are zeros and therefore we have that the sequence $[s_{1},s_{2},\ldots,s_{2^{m-1}-1}]$ is the shortened half de Bruijn sequence obtained from ${\cal S}$ (starts with $m-1$ zeros). Define the sequence
$B\triangleq(\overbrace{0,0,\ldots,0}^{2^{t}~{}\text{copies}},\overbrace{s_{1},s_{2},\ldots,s_{2^{m-1}-1},\ldots,s_{1},s_{2},\ldots,s_{2^{m-1}-1}}^{2^{t}~{}\text{copies~{}of}~{}s_{1},s_{2},\ldots,s_{2^{m-1}-1}})$
of length $2^{t}+(2^{m-1}-1)2^{t}=2^{m+t-1}$ and denote $B=(b_{1},b_{2},\ldots,b_{2^{m+t-1}})$.
We define a new array ${\cal B}$ in a new code $\mathbb{C}^{\prime}$. Let the $i$-th column of ${\cal B}$ be defined by, ${\cal B}_{i}={\bf D}^{-1}_{b_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of ${\cal A}$ and $1\leq i\leq 2^{m+t-1}$. The codeword ${\cal B}$ is a $2^{k+1}\times 2^{m+t-1}$ doubly-periodic array and hence it has $2^{k+m+t}$ windows of size $(n+1)\times m$. Since $\Delta 2^{k+m+t}=2^{m+nm}=2^{(n+1)m}$ (because $\Delta 2^{k+t}=2^{nm}$), it follows that to complete the proof it is sufficient to show that all the $(n+1)\times m$ windows in all the codewords of $\mathbb{C}^{\prime}$ are distinct.
The rest is the proof is very similar to the one of Theorem 12 and hence it is omitted. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k+1},2^{m+t-1};n+1,m)$-DBAC of size $\Delta$. ∎
The construction of Theorem 15 yields DBACs with new parameters in a similar way to the constructions in Theorems 12 and 13.
Finally, we should note that in some of the recursive constructions also columns of length $2^{k}$, but with different periods can be used. We did not use them in the constructions to avoid more complexity.
V Joining Cycles in the Factor Graph
In this section, we provide a construction for $(2^{k},2^{t};n,m)$-DBACs from $(2^{k},2^{t^{\prime}};n,m)$-DBACs, where $t\geq t^{\prime}$. This construction is based on joining cycles of the previous constructed DBACs, when the codewords are considered as cycles in the graph $G_{{\cal F},m}$. The construction starts with a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where all the columns of the codewords are from ${\cal F}$, a perfect factor PF$(n,k)$ and hence they are cycles in $G_{{\cal F},m}$. The codewords should be ordered in pairs such that each pair of codewords has a pair of companion states, i.e., if $({\cal C}_{1},{\cal C}_{2})$ is such a pair of codewords, then there are two companion vertices, one on ${\cal C}_{1}$ and a second on ${\cal C}_{2}$. Now, we join each pair by their companion vertices using the merge-or-split method. The obtained code is a $(2^{k},2^{t+1};n,m)$-DBAC of size $\Delta/2$. This process can be continue and applied on the new generated DBAC. We should note that if the sum of shifts in one codeword is congruent to 0 modulo $2^{k}$ and in the second codeword the sum is also congruent to 0 modulo $2^{k}$, then in the merged cycles (codewords) the sum of shifts is also congruent to 0 modulo $2^{k}$. Therefore, merging two cycles which are represented by two doubly-periodic arrays, using the merge-or-split method, yields a new cycle which can be represented by a doubly-periodic array.
We will demonstrate the idea with two examples based on the codes obtained in Theorems 8 and 9. The first example will be comprehensive while the second example will be only partial.
The first example of such a method is to start with the cycles of a $(2^{k},2^{t+1};n,2^{t})$-DBAC $\mathbb{C}$ obtained from the construction of Theorem 9 with ${\cal F}$ as the PF$(n,k)$. This code has $2^{n2^{t}-k-t-1}$ codewords. Our goal is to order all the codewords in a list (cyclic as all the lists which will be generated in the two examples in this section) such that each two consecutive codewords in this list have a pair of companion vertices and hence they can be merge together using the merge-or-split method. Since this list has $2^{n2^{t}-k-t-1}$ codewords, it will be possible to merge any sequence of $2^{s}$ consecutive codewords, where $10\leq s\leq n2^{t}-k-t-$ 1 and obtain a $(2^{k},2^{t+1+s};n,2^{t})$-DBAC.
We start by any ordering the cycles of ${\cal F}$
$X_{0},X_{1},\ldots,X_{2^{n-k-1}-1},X_{2^{n-k-1}},\ldots,X_{2^{n-k}-1},$
where $X_{2^{n-k-1}+i}=\bar{X}_{i}$, where $10\leq i\leq 2^{n-k-1}-$ 1. We apply the following simple mapping
$\varphi:{\cal F}\longrightarrow\mathbb{F}_{2}$
where $\varphi(X_{i})=0$ if $10\leq i\leq 2^{n-k-1}-$ 1 and $\varphi(X_{i})=1$ if $12^{n-k-1}\leq i\leq 2^{n-k}-$ 1. The mapping $\varphi$ induces a mapping from the codewords of $\mathbb{C}$ into the set of self-dual sequences of length $2^{t+1}$. Such a mapping maps each codeword of $\mathbb{C}$ into exactly one self-dual sequence of length $2^{t+1}$.
Lemma 13
The number of codewords of $\mathbb{C}$ which are mapped into a given self-dual sequence ${\cal S}$ of length $2^{t+1}$ is $2^{(n-k-1)2^{t}}2^{k(2^{t}-1)}$.
Proof:
Given a self-dual sequence ${\cal S}$ of length $2^{t+1}$, it has exactly $2^{t}$ coordinates with zeros. On each of these coordinates we can choose any one of the first $2^{n-k-1}$ cycles from $\{X_{0},X_{1},\ldots,X_{2^{n-k-1}-1}\}$ for a total of $2^{(n-k-1)2^{t}}$ distinct choices. Fixing any coordinate with a zero and taking its cycle in its zero shift, in each one of the other $12^{t}-$ 1 coordinates with zeros, we can use any one of the $2^{k}$ possible shifts of the cycle compared to its zero state for a total of $2^{k(2^{t}-1)}$ distinct choices. Thus, the claim of the lemma is proved. ∎
It was proved in [14] that the self-dual sequences of length $2^{t+1}$, $t\geq 2$, can be ordered in a way that each two consecutive sequences differ in exactly two positions $\ell$ and $\ell+2^{t}$. This ordering will be very useful in obtaining the final list with all the codewords in $\mathbb{C}$. All the codewords which are mapped into the same self-dual sequence will be partitioned again into equivalence classes to simplify the understanding of the ordering. Consider a self-dual sequence ${\cal S}$ of length $2^{t+1}$ and its $2^{t}$ coordinates which are zeros. One of them is fixed as the first one and let
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$
be the cycles of ${\cal F}$, in these $2^{t}$ coordinates given in any order, such that $10\leq i_{j}\leq 2^{n-k-1}-$ 1. There are $2^{k(2^{t}-1)}$ codewords of this type for ${\cal S}$ (having $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ in these $2^{t}$ coordinates) associated with the shifts of these cycles compared to their zero states, where $X_{i_{1}}$ will be taken in its zero state.
The next step is to order these $2^{k(2^{t}-1)}$ codewords in a way that each two consecutive codewords have a pair of companion vertices. For this purpose we consider the set of ordered shifts of $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$, i.e., let
${\cal Q}_{1}\triangleq\{(0,j_{2},j_{3},\ldots,j_{2^{t}})~{}:~{}0\leq j_{\ell}\leq 2^{k}-1,~{}2\leq\ell\leq 2^{t}\},$
where $(0,j_{2},j_{3},\ldots,j_{2^{t}})$ with $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ is associated with the following shifts of these cycles of ${\cal F}$ in a codeword ${\cal C}\in\mathbb{C}$,
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}}),$
where these entries are usually not consecutive in the codeword. Hence, there exists a one-to-one mapping between the elements in the set ${\cal Q}_{1}$ and the number of codewords represented by
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$
for the self-dual sequence ${\cal S}$.
Our next step in combining the codewords is to order the $2^{k(2^{t}-1)}$ elements of ${\cal Q}_{1}$ in a cyclic Gray code, i.e., that each two elements will differ in exactly one position. Such a cyclic Gray code is well-known [38] and is also very easy to construct. The following lemma can be readily verified.
Lemma 14
If $(0,j_{2},j_{3},\ldots,j_{2^{t}})\in{\cal Q}_{1}$ and $(0,j^{\prime}_{2},j^{\prime}_{3},\ldots,j^{\prime}_{2^{t}})\in{\cal Q}_{1}$ differ in exactly one coordinate, then their associated codewords in the same equivalence class
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}}),$
and
$(X_{i_{1}},{\bf E}^{j^{\prime}_{2}}X_{i_{2}},\ldots,{\bf E}^{j^{\prime}_{2^{t}}}X_{i_{2^{t}}}).$
can be combined into one codeword by the merge-or-split method, i.e., they have a pair of companion vertices.
Proof:
The two codewords in $Q_{1}$ differ in exactly one coordinate implies that the associated two codewords of $\mathbb{C}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, differ in exactly two coordinates separated by $12^{t}-$ 1 coordinates in which both ${\cal C}_{1}$ and ${\cal C}_{2}$ are identical. The two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ associated with these $12^{t}-$ 1 coordinates and the coordinate which follows them differ only in the last column. This implies that these two $2^{k}\times 2^{t}$ matrices of ${\cal C}_{1}$ and ${\cal C}_{2}$ are companion vertices. ∎
The cyclic Gray code of the elements of the set ${\cal Q}_{1}$ and Lemma 14 implies that we can form a cyclic Gray code from the codewords of each equivalence class.
The next step is to order the codewords from different equivalence classes associated with the same self-dual sequence ${\cal S}$. For this purpose we define another Gray code on the set of elements $Q_{2}$ defined by
${\cal Q}_{2}\triangleq\{(i_{1},i_{2},i_{3},\ldots,i_{2^{t}})~{}:~{}0\leq i_{\ell}\leq 2^{n-k-1}-1,~{}1\leq\ell\leq 2^{t}\}.$
The element $(i_{1},i_{2},i_{3},\ldots,i_{2^{t}})\in Q_{2}$ is associated with the equivalence class
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}}).$
It is easy to verify again where we can find a pair of companion vertices in two distinct equivalence classes associated with the same self-dual sequence ${\cal S}$, which differ in exactly one coordinate of $Q_{2}$.
Lemma 15
If $(i_{i},i_{2},i_{3},\ldots,i_{2^{t}})\in{\cal Q}_{2}$ and $(i^{\prime}_{1},i^{\prime}_{2},i^{\prime}_{3},\ldots,i^{\prime}_{2^{t}})\in{\cal Q}_{1}$ differ in exactly one coordinate, then their associated codewords (where all the cycles are the same are also in the same shift) in the associated equivalence classes
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}}),$
and
$(X_{i^{\prime}_{1}},X_{i^{\prime}_{2}},\ldots,X_{i^{\prime}_{2^{t}}}),$
can be combined into one codeword by the merge-or-split method, i.e., they have a pair of companion vertices.
Proof:
These two codewords differs in two coordinates as is Lemma 14 and hence have a pair of companion states. ∎
We continue to consider two self-dual sequences of length $2^{t+1}$, ${\cal S}_{1}$ and ${\cal S}_{2}$ which differ in exactly two coordinates $i$ and $i+2^{t}$.
Lemma 16
Any two self-dual sequences of length $2^{t+1}$, ${\cal S}_{1}$ and ${\cal S}_{2}$ which differ in exactly two coordinates have in their associated codewords a pair of codewords whose vertices have a pair of companion vertices.
Proof:
As before we can use two associated codewords which differ in their cycles only in these two coordinates. Again, we can use all the cycles in their zero states and hence the two associated codewords will have a pair of companion vertices. ∎
The last important step is to use Lemmas 1415, and 16 to combine all the cyclic lists into one cyclic list which contains all the codewords and each two consecutive codewords have a pair of companion vertices.
Theorem 16
All the codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC $\mathbb{C}$ obtained in the construction of Theorem 9 can be ordered in a cyclic list in which each two consecutive codewords have a pair of companion vertices.
Proof:
We start by merging together two cyclic lists associated with the same self-dual sequence ${\cal S}$ and two different equivalence classes which are the first two in the list of equivalence classes associated with $Q_{2}$,
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}}),$
and
$(X_{i^{\prime}_{1}},X_{i^{\prime}_{2}},\ldots,X_{i^{\prime}_{2^{t}}}),$
which differ in exactly one coordinate. Assume further that the first element in the Gray code list of $Q_{1}$ is
$(0,j_{2},j_{3},\ldots,j_{2^{t}})$
and the last element is
$(0,j^{\prime}_{2},j^{\prime}_{3},\ldots,j^{\prime}_{2^{t}})~{}.$
The combined list starts with
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}}),$
and continue with the elements of the same equivalence class until it reaches
$(X_{i_{1}},{\bf E}^{j^{\prime}_{2}}X_{i_{2}},\ldots,{\bf E}^{j^{\prime}_{2^{t}}}X_{i_{2^{t}}}),$
The next element in the combined list will be
$(X_{i^{\prime}_{1}},{\bf E}^{j^{\prime}_{2}}X_{i^{\prime}_{2}},\ldots,{\bf E}^{j^{\prime}_{2^{t}}}X_{i^{\prime}_{2^{t}}}).$
These last two elements differ in exactly one coordinate and hence they are associated with a pair of companion vertices. We continue with this new equivalence class in reverse order until we reach the element
$(X_{i^{\prime}_{1}},{\bf E}^{j_{2}}X_{i^{\prime}_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i^{\prime}_{2^{t}}}),$
which differ in exactly one coordinate from the first element
$(X_{i_{1}},{\bf E}^{j^{\prime}_{2}}X_{i_{2}},\ldots,{\bf E}^{j^{\prime}_{2^{t}}}X_{i_{2^{t}}}).$
We can continue in the same manner with the third equivalence class in the list in its usual order and the fourth equivalence class in reverse order and so on until all the equivalence classes associated with ${\cal S}$ are merged together.
Once all the codewords of the different equivalence classes for each self-dual sequence were merged together into one cyclic list in which each two consecutive codewords have a pair of companion vertices we have to combine the different codewords of distinct self-dual sequences into one long cyclic lict in which each two consecutive codewords have a pair of companion vertices. Let $S_{1}$ and $S_{2}$ be the first two self-dual sequences in the list. By the definition, ${\cal S}_{1}$ and ${\cal S}_{2}$ differ in exactly two coordinates $\ell$ and $\ell+2^{t}$. We take for these two self-dual sequences two equivalence classes that differ only in the coordinates in which ${\cal S}_{1}$ and ${\cal S}_{2}$ differ and merge the two list using the same idea as the one in merging the lists of two different equivalence classed for the same self-dual sequence.
Thus, the outcome is a cyclic list with all the codewords of $\mathbb{C}$, in which each two consecutive codewords have a pair of companion vertices. ∎
Corollary 4
If $k$ and $n$ are integers such that $k<n<2^{k}$, then there exist a $(2^{k},2^{t+1+s};n,2^{t})$-DBAC, where $10\leq s\leq n2^{t}-k-t-$ 1, of size $2^{n2^{t}-k-t-1-s}$.
An example for the list of codewords of a $(2^{k},2^{t+1};n,2^{t})$-DBAC such that each two codewords have a pair of companion state is too large and hence we omit such an example and leave it for the interested reader.
A second example of the method is to start with the codewords of a $(2^{k},2^{t};n,2^{t}-1)$-DBAC $\mathbb{C}$ obtained from the construction of Theorem 8 with ${\cal F}$ as the PF$(n,k)$. This code has $2^{nm-k-t}$ codewords, where $m=2^{t}-1$ and $t\geq k$. We start with one partition of these codewords into equivalence classes (disjoint sets) as follows. Each codeword ${\cal C}\in\mathbb{C}$ is represented as a word of length $2^{t}$ as follows:
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}}),$
where this representation is the one among the $2^{t}$ cyclic shifts of ${\cal C}$ with the smallest lexicographic value for $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ (although this lexicographic value will have no influence on the construction).
We define an equivalence relation on all the codewords of $\mathbb{C}$. The codeword ${\cal C}$ will be in the equivalence class
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}}),$
Each $X_{i_{j}}$, except for the last one, can be chosen in $2^{n-k}$ distinct ways for a total of $2^{(n-k)(2^{t}-1)}$ distinct choices. the last one is determined by the equation $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$. Since each one of the $2^{t}$ cycles from ${\cal F}$ in the equivalence class
$(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$
can be chosen as the first one, it follows that the number of equivalence classes in this equivalence relation is
$2^{(n-k)(2^{t}-1)}/2^{t}=2^{(n-k)(2^{t}-1)-t}~{}.$
Merging the codewords in each equivalence class can be made exactly as in the first example. The first sequence is taken in its zero shift, The next $22^{t}-$ 2 sequences can be taken each one in any of its $2^{k}$ distinct cyclic shifts. The shift of the last sequence it determined in a way that the sum of all the shifts is congruent to 0 modulo $2^{k}$. Two codewords of an equivalence class which differ in exactly two consecutive shifts have a pair of companion states since the number of columns of a state is $12^{t}-$ 1 and these two codewords can be horizontally shifted in a way that they share the first $22^{t}-$ 2 columns and their shifts. For this purpose we consider the set $Q_{3}$ whose elements define the shifts of the sequences from ${\cal F}$. The set $Q_{3}$ is defined as follows.
${\cal Q}_{3}\triangleq\{(0,j_{2},j_{3},\ldots,j_{2^{t}})~{}:~{}0\leq j_{\ell}\leq 2^{k}-1,~{}2\leq\ell\leq 2^{t},~{}\sum_{\ell=2}^{2^{t}}j_{\ell}\equiv 0~{}({\mbox{mod}}~{}2^{k})\},$
where the element $(0,j_{2},j_{3},\ldots,j_{2^{t}})\in Q_{3}$ for the equivalence class $(X_{i_{1}},X_{i_{2}},\ldots,X_{i_{2^{t}}})$ is associated with the codeword
$(X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{2^{t}}}X_{i_{2^{t}}}).$
Now, we should order the elements of $Q_{3}$ in a cyclic list such that each two elements in the list differ in exactly two consecutive elements. Such a list implies that each number of consecutive elements in the least can be merged together via the merge-or-split method. To have a longer cyclic list which includes all the codewords of $\mathbb{C}$ we have to order the equivalence classes in another cyclic list, where each two consecutive equivalence classes in the list differ in exactly two consecutive elements. Note, that these equivalence classes are over an alphabet of size $2^{n-k}$, the number of sequences in ${\cal F}$. Two codewords from consecutive elements in such a list which are associated with the same shifts from $Q_{3}$ differ in exactly two coordinates and hence have a pair of companion states. We do not have a general construction for such a list and we leave it as a problem for future research. When the number of sequence in ${\cal F}$ is two, i.e., $n-k=1$, it is relatively easy to generate such a list. The list consists of all the cyclic sequences of length $2^{t}$ and odd weight. We should order these elements in a list in which each two consecutive elements differ in exactly two coordinates. For this purpose we consider the list of self-dual sequences of length $2^{t+1}$. If $Y_{i}$ and $Y_{i+1}$ are two consecutive elements in this list, then $Y_{i}$ and $Y_{i+1}$ differ in two positions separated by $12^{t}-$ 1 other positions. ${\bf D}Y_{i}$ and ${\bf D}Y_{i+1}$ have odd weight and they differ in exactly two adjacent positions. Thus, the existence of a list for the self-dual sequences of length $2^{t+1}$ implies the list of sequences of odd weight and length $2^{t}$.
Example 3
If $n-k=1$, then there are two sequences in a PF$(n,k)$, $0001$ and $0111$ and these two sequences are the columns of all the codewords in $\mathbb{C}$. When $t=2$ the equivalence classes are also $0001$ and $0111$. The ordered list of these two equivalence classes is trivial. A list with all the elements of $Q_{3}$ is given as follows (the elements of $Q_{3}$ are presented as column vectors).
$\begin{array}[]{c}0000000000000000\\ 0110033332211022\\ 0301301323121202\\ 0033110233112200\end{array}$
The code $\mathbb{C}$ has 32 codewords. The 16 codewords of the equivalence class $0001$ are ordered in the list as follows.
$\left[\begin{array}[]{c}0000\\ 0001\\ 0001\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0010\\ 0001\\ 0101\\ 1001\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0000\\ 0101\\ 1011\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0000\\ 0011\\ 1101\end{array}\right],~{}\left[\begin{array}[]{c}0011\\ 0001\\ 0001\\ 1100\end{array}\right],\left[\begin{array}[]{c}0101\\ 0001\\ 0001\\ 1010\end{array}\right],~{}\left[\begin{array}[]{c}0100\\ 0001\\ 0011\\ 1001\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0001\\ 0000\\ 1001\end{array}\right],$
$\left[\begin{array}[]{c}0101\\ 0010\\ 0001\\ 1001\end{array}\right],~{}\left[\begin{array}[]{c}0011\\ 0100\\ 0001\\ 1001\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0101\\ 0011\\ 1000\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0011\\ 0101\\ 1000\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0001\\ 0110\\ 1001\end{array}\right],\left[\begin{array}[]{c}0001\\ 0011\\ 0000\\ 1101\end{array}\right],~{}\left[\begin{array}[]{c}0001\\ 0101\\ 0000\\ 1011\end{array}\right],~{}\left[\begin{array}[]{c}0000\\ 0111\\ 0001\\ 1001\end{array}\right],$
The 16 codewords of the equivalence class $0111$ are ordered in the list as follows.
$\left[\begin{array}[]{c}0000\\ 0111\\ 0111\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0110\\ 0101\\ 0111\\ 1011\end{array}\right],~{}\left[\begin{array}[]{c}0101\\ 0110\\ 0111\\ 1011\end{array}\right],~{}\left[\begin{array}[]{c}0011\\ 0110\\ 0111\\ 1101\end{array}\right],~{}\left[\begin{array}[]{c}0011\\ 0101\\ 0111\\ 1110\end{array}\right],~{}\left[\begin{array}[]{c}0101\\ 0011\\ 0111\\ 1110\end{array}\right],~{}\left[\begin{array}[]{c}0110\\ 0011\\ 0111\\ 1101\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0001\\ 0110\\ 1111\end{array}\right],$
$\left[\begin{array}[]{c}0111\\ 0010\\ 0101\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0100\\ 0011\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0111\\ 0011\\ 1100\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0111\\ 0101\\ 1010\end{array}\right],~{}\left[\begin{array}[]{c}0111\\ 0111\\ 0110\\ 1001\end{array}\right],~{}\left[\begin{array}[]{c}0011\\ 0111\\ 0100\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0101\\ 0111\\ 0010\\ 1111\end{array}\right],~{}\left[\begin{array}[]{c}0110\\ 0111\\ 0001\\ 1111\end{array}\right].$
Now, it is easy to verify that any set of consecutive codewords can be merged together. Hence, we can obtain a $(4,2^{2+i};3,3)$-DBAC for each $i$, $0\leq i\leq 5$,
VI Conclusion
We have defined the concept of de Bruijn array codes which generalizes the concept of a de Bruijn array. The main interesting question is whether the necessary conditions of Lemma 1 are also sufficient. In other words, given four integers, $k$, $t$, $n$, $m$, such that $\Delta=2^{nm-k-t}$, does there exists a $(2^{k},2^{t};n,m)$-DBAC with $\Delta$ codewords? We conjecture that such a code exists, i.e., the necessary conditions are sufficient.
If $\Delta=1$ the problem was completely solved by Paterson [42]. The direct constructions together with the merge-or-split method to merge codewords and the recursive construction solve some families of parameters. First, these constructions were designed for the case where a PF$(n,k)$ exists, i.e., when $k\leq n<2^{k}$, i.e., $n<2^{k}\leq 2^{n}$, but by the transposing the codeword and applying the recursive constructions other sets of parameters can be obtained.
To prove that the necessary conditions are indeed sufficient more constructions are required. These are left for future research. Indeed, the span of parameters for different families are so wide that it will probably take some time before this will ultimately by solved. Similarly to de Bruijn array codes, we can define pseudo-random array codes, a generalization for pseudo-random arrays. While for de Bruijn array codes most of the constructions have a combinatorial flavor, for pseudo-random array codes all the constructions will have an algebraic flavor. These arrays were considered in [11] and a comprehensive work on them will be presented in the follow-up paper [5].
The description of joining codewords from the codes obtained from the construction of Theorem 8 leads to two open questions associated with ordering of codewords of a given code $\mathbb{C}$.
Let $Q_{4}$ be the set defined by
${\cal Q}_{4}\triangleq\{(i_{1},i_{2},i_{3},\ldots,i_{2^{t}})~{}:~{}0\leq i_{\ell}\leq 2^{n-k}-1,~{}1\leq\ell\leq 2^{t},~{}\sum_{\ell=1}^{2^{t}}i_{\ell}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})\}.$
  1. 1.
    Order the elements of $Q_{3}$ in a cyclic list such that any two consecutive elements in the list differ in exactly two adjacent coordinates.
  2. 2.
    Order the elements of $Q_{4}$ in a cyclic list such that any two consecutive elements in the list differ in exactly two adjacent coordinates.
Acknowledgement
Part of this work have been presented at the IEEE International Symposium on Information Theory, Athens, Greece,, July 2024. This research was supported in part by the Israeli Science Foundation grant no. 222/19.